### v2

# When is Multicalibration Post-Processing Necessary?

Abstract
Calibration is a well-studied property of predictors which guarantees meaningful uncertainty estimates. Multicalibration is a related notion — originating in algorithmic fairness — which requires predictors to be simultaneously calibrated over a potentially complex and overlapping collection of protected subpopulations (such as groups defined by ethnicity, race, or income). We conduct the first comprehensive study evaluating the usefulness of multicalibration post-processing across a broad set of tabular, image, and language datasets for models spanning from simple decision trees to 90 million parameter fine-tuned LLMs. Our findings can be summarized as follows: (1) models which are calibrated out of the box tend to be relatively multicalibrated without any additional post-processing; (2) multicalibration post-processing can help inherently uncalibrated models and large vision and language models; and (3) traditional calibration measures may sometimes provide multicalibration implicitly. More generally, we distill many independent observations which may be useful for practical and effective applications of multicalibration post-processing in real-world contexts. We also release a python package implementing multicalibration algorithms, available via ‘pip install multicalibration’.
* Equal contribution.
1 Introduction
A popular approach to ensuring that probabilistic predictions from machine learning algorithms are meaningful is model calibration. Intuitively, calibration requires that amongst all samples given score $p\in[0,1]$ by an ML algorithm, exactly a $p$-fraction of those samples have positive label. Calibration ensures that a predictor has an accurate estimate of its own predictive uncertainty, and is a fundamental requirement in applications where probabilities may be taken into account for high-stake decisions such as disease diagnosis (Dahabreh et al., , 2017) or credit/lending decisions (Bequé et al., , 2017). Miscalibration can result in undesirable downstream consequences when probabilistic predictions are thresholded into decisions: if a predictor has high calibration error in disease diagnosis, for example, the individuals assigned lower predicted probabilities may be unfairly denied treatment. Calibration has a long history in the machine learning community (Platt et al., , 1999; Niculescu-Mizil and Caruana, , 2005; Guo et al., , 2017; Minderer et al., , 2021), but was arguably first introduced in fairness contexts by Cleary, (1968). More recently, it has appeared in the algorithmic fairness community via the seminal works of Chouldechova, (2017); Kleinberg et al., (2017).
Although calibration ensures meaningful uncertainty estimates aggregated over the entire population, it does not preclude potential discrimination at the level of groups of individuals: a model may be well calibrated overall but systematically underestimate the risk or qualification probability on historically underrepresented subsets of individuals. For example, Obermeyer et al., (2019) show differing calibration error rates across groups defined by race for prediction in high-risk patient care management systems. As pointed out by Obermeyer et al., (2019), in the downstream task of patient intervention based on thresholds over probabilistic predictions, this can inadvertently lead to differing rates of healthcare access based on group membership.
To combat these issues, the notion of multicalibration was proposed as a refinement of standard calibration (Hébert-Johnson et al., , 2018). Multicalibration requires that a model be simultaneously calibrated on an entire collection of (efficiently) identifiable and potentially overlapping subgroups of the data distribution. A plethora of recent theoretical work has studied and utilized multicalibration to obtain interesting and important guarantees in algorithmic fairness (Gopalan et al., 2022b, ; Gopalan et al., 2022c, ; Devic et al., , 2024; Dwork et al., , 2021; Shabat et al., , 2020; Jung et al., , 2021; Bastani et al., , 2022), learning theory (Gopalan et al., 2022a, ; Gollakota et al., , 2024; Gopalan et al., , 2023), and cryptography (Dwork et al., , 2023). Desirable consequences of multicalibrated predictors abound: multicalibration can provide provable guarantees on the transferability of a model’s predictions to different loss functions (omniprediction, Gopalan et al., 2022a ), the ability of a model to do meaningful conformal prediction (Jung et al., , 2023), and universal adaptability or domain adaptation (Kim et al., , 2022).
Although there is a host of theoretical results surrounding multicalibration and related notions, there is little systematic empirical study of the latent multicalibration error of popular machine learning models, the effectiveness of multicalibration post-processing algorithms, or even best practices for practitioners who wish to apply ideas and algorithms from the multicalibration literature. In particular, theoretical results are often concerned with multicalibration towards subgroups defined by potentially infinite hypothesis classes (Hébert-Johnson et al., , 2018; Haghtalab et al., , 2023). In contrast, fairness practitioners may prioritize the equitable performance of a model over a finite number of protected subgroups of interest. These groups are typically defined by attributes and meta-data such as race, sex, ethnicity, etc. (Chen et al., , 2023) which are normatively deemed as important. Furthermore, most existing works applying multicalibration in practical settings only focus on one-off datasets or examples, and do not validate the algorithm(s) across a variety of datasets and models or with realistic finite sample restrictions (Liu et al., , 2019; Barda et al., , 2020; La Cava et al., , 2023).
To address these, we consider a “realistic” setup where a practitioner only has a finite amount of data, and must choose how to partition this data between learning and post-processing in order to achieve a suitable accuracy and multicalibration error rate over a finite set of subgroups. This allows us to investigate many important questions pertaining to the practical usage of multicalibration concepts and algorithms, which, to the best of our knowledge, have not been systematically considered by the theoretical or practical communities. For example, we use this setup to investigate the effectiveness of multicalibration post-processing algorithms and hyperparameter choices, as well as the latent multicalibration properties of popular machine learning models at a large scale.
More broadly, we initiate a systematic empirical study of multicalibration with the goal of answering two salient questions:
Question 1. In practice, how often and for what machine learning models is multicalibration an expected consequence of empirical risk minimization?
Question 2. Conversely, when must additional steps be taken to multicalibrate models, how difficult is this to do in practice, and what steps can be taken to make this easier?
The conventional wisdom is that multicalibration is something that is not naturally achieved by ML algorithms—this is precisely why many in the community have focused on creating post-processing algorithms which do achieve it (see e.g. Hébert-Johnson et al., (2018); Gopalan et al., 2022b , and Section 1.2). However, recent theoretical results suggest that multicalibration may in fact be an inevitable consequence of certain empirical risk minimization (ERM) methods with proper losses (Błasiok et al., , 2023; Liu et al., , 2019). This apparent conflict between conventional wisdom and recent results has not been tested in practice. We propose studying Question 1 since we believe that the current state of multicalibration in ML models should be systematically studied to better understand the implications for modern learning setups involving large models and fine-tuning. Question 2 is complementary and focused on investigating the effectiveness of current multicalibration algorithms on real datasets and illuminating challenges which can guide the development of future algorithms. A partial answer to one or both of these questions could help practitioners concerned about fairness understand when they should or should not expect multicalibration algorithms to help.
Refer to captionRefer to captionRefer to caption
Figure 1: Test accuracy vs. maximum group-wise calibration error (smECE) averaged over five train/validation splits for simple neural networks (MLPs) trained on Credit Default, MEPS, and ACS Income. Each point corresponds to the performance of the multicalibration post-processing algorithm HKRR (Hébert-Johnson et al., , 2018) or HJZ (Haghtalab et al., , 2023) with a different choice of hyperparameters. Standard empirical risk minimization (ERM) for MLPs achieves nearly optimal accuracy and multicalibration error. Full hyperparameter plots for each base dataset are in Appendix F.
1.1 Our Contributions
We conduct a large-scale evaluation of multicalibration methods, comparing three families of methods: (1) standard ERM, (2) ERM followed by a classical recalibration method (e.g. Platt scaling), and (3) ERM followed by an explicit multicalibration algorithm (e.g. that of Hébert-Johnson et al., (2018)).
We find that in practice, this comparison is surprisingly subtle: multicalibration algorithms do not always improve worst group calibration error (relative to the ERM baseline), for example. From the results of our extensive experiments on tabular, vision, and language tasks (involving running multicalibration algorithms more than 45K times), we extract a number of observations clarifying the utility of multicalibration algorithms. Most significantly, we find:
  1. 1.
    ERM alone is often a strong baseline, and can often be remarkably multicalibrated without further post-processing. In particular, on tabular datasets, multicalibration post-processing does not improve upon worst group calibration error of ERM for simple NNs.
  2. 2.
    Multicalibration algorithms are very sensitive to hyperparameter choices, and can require large parameter sweeps to avoid overfitting. Furthermore, these algorithms tend to be most effective in regimes with large amounts of available data, such as image and language datasets.
  3. 3.
    Traditional calibration methods such as Platt scaling or isotonic regression can sometimes give nearly the same performance as multicalibration algorithms, and are hyperparameter-free. Furthermore, compared to multicalibration post-processing, they are extremely computationally efficient.
We also present numerous practical takeaways for users of multicalibration algorithms, which are not apparent from the existing theoretical literature, but are crucial considerations in practice. We believe that our investigations will not only broaden the practical applicability of multicalibration as a concept and algorithm, but also provide valuable information to the theoretical community as to what barriers multicalibration faces in practice. To both of these ends, all code used in our experiments is publicly accessible, and we also release a python package implementing two multicalibration algorithms which we make available via ‘pip install multicalibration’.  [^1] Code for our experiments is available at https://github.com/dutchhansen/empirical-multicalibration, while code for the python package is available at https://github.com/sid-devic/multicalibration.
Organization. In Section 1.2, we begin with a brief review of related theoretical and experimental work in the multicalibration literature. We then detail our key experimental design choices in Section 2, before discussing our results on tabular data in Section 3. We extend our results to more complex image and language datasets in Section 4. In Section 5, we investigate to what extent multicalibration post-processing can re-use data from model training. Finally, we conclude with limitations of our experiments as well as practical takeaways for practitioners of fair machine learning in Section 6.
1.2 Related Works: Theory and Practice
The theory of multicalibration is rife with theoretical results investigating the sample complexity (Shabat et al., , 2020), learnability, and computational efficiency of multicalibrated predictors. Hébert-Johnson et al., (2018) initiated this study by showing that achieving multicalibration over a hypothesis class $\mathcal{C}$ defining protected subgroups requires access to a weak agnostic learner for that class (Shalev-Shwartz and Ben-David, , 2014). From a fairness perspective, however, we are oftentimes—but not always (Sahlgren and Laitinen, , 2020)—interested in subgroups defined by features or metadata, rather than a generic (and potentially infinite) hypothesis class. Subgroups in practical applications of algorithmic fairness are often given as input to the machine learning algorithm and intrinsic to a particular dataset of interest.
There is reason to believe that empirical risk minimization (ERM) on neural networks and other machine learning models may result in multicalibrated predictors. In recent works, Błasiok et al., (2024, 2023) prove that loss minimization with neural networks may yield multicalibrated predictors. Their proofs, however, may not be directly applicable to practice as they rely on an idealized optimization procedure (we provide further discussion of the relation between our works in Section A.1). Nonetheless, both works echo a relationship between ERM and multicalibration also articulated in Liu et al., (2019), who show that group-wise calibration may be an inevitable consequence of well-performing models.
Although these results describe and prove links between ERM and multicalibration in theory, we systematically evaluate when this link holds in practice across a broad range of models. To the best of our knowledge, only La Cava et al., (2022); Barda et al., (2021) consider issues when applying multicalibration in practice. Both works are limited to small models or only run experiments with one or two datasets. Pfohl et al., (2022) measure subgroup calibration, but do not discuss it at length. Although the focus in Haghtalab et al., (2023) is mainly theoretical, the authors also run experiments for both their proposed multicalibration algorithms and additional algorithms present in the literature. However, they do not compare the performance of such algorithms against a baseline NN or tree-based method trained on an identical amount of data, and also do not investigate how much data should be used for training the base model vs. used for multicalibration post-processing. We believe such a comparison is important given that ERM may have latent multicalibration properties, and is by far the most used learning algorithm in practice.
In recent work, Detommaso et al., (2024) utilize multicalibration as a tool to improve the overall uncertainty and confidence calibration of language models but, to our knowledge, do not focus on or report fairness towards protected subgroups. Like us, they point out various issues with the standard multicalibration algorithm, which they address with early stopping and adaptive binning. We instead perform a large hyperparameter sweep which effectively implements an early stopping mechanism. We discuss this further in Section A.2. Nonetheless, our results for large models are complementary to those of Detommaso et al., (2024): both works demonstrate that (1) standard multicalibration can at times be difficult to get working in practice; and (2) ideas from the theoretical multicalibration literature can have impact at the scale of large models. We provide additional discussion of related works in model calibration and subgroup robustness in Appendix A.
2 Preliminaries
We work in the binary classification setting with a domain $\mathcal{X}$ and binary label set $\mathcal{Y}=\{0,1\}$, and assume data is drawn from a distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$. We consider arbitrary risk predictors $f:\mathcal{X}\rightarrow\Delta(\mathcal{Y})$, which return probability distributions over the binary label space. We will measure the calibration of $f$ on a dataset $S\in(\mathcal{X}\times\mathcal{Y})^{n}$ with the binned variant of the well-known and standard Expected Calibration Error, which we refer to as ECE (Guo et al., , 2017). Throughout, we measure ECE with 10 bins of equal width $0.1$.
Recent work has questioned ECE as a calibration measure, due to consistency and continuity issues that come with relying on a fixed bin width (Błasiok et al., , 2023). To address these, we also report calibration as measured by smoothed ECE (smECE) (Błasiok and Nakkiran, , 2023), which (1) can be roughly thought of as the ECE after applying a suitable kernel-smoothing to the predictions, and (2) satisfies desirable continuity and consistency guarantees (Błasiok et al., , 2023; Błasiok and Nakkiran, , 2023). Importantly, unlike binned ECE, there are no hyperparameters associated with measuring the smoothed calibration error. A full description of smECE is beyond the scope of our work—we refer the interested reader to Błasiok and Nakkiran, (2023).
Multicalibration requires that a predictor have not only small calibration error overall, but also when restricted to marginal subgroup distributions of the data. In particular, we assume that there is a (finite) collection of groups $G=\{g_{1},g_{2},\dots\}$, where $g_{i}\subseteq\mathcal{X}$. We operationalize measuring multicalibration by reporting the maximum calibration error over a given collection of subgroups $G$ [^2] Multicalibration was introduced before smECE, and was designed to reduce a bucketed group-wise calibration error (similar to ECE). Therefore, our investigations concerning smECE are of a purely empirical character. Taking the max avoids fairness concerns associated with the (weighted) mean of groups of varying size and/or degree of overlap. Note that the subgroup collection $G$ is context and dataset dependent, and that the groups within $G$ may be overlapping, capturing desirable intersectionality notions (Ovalle et al., , 2023).
2.1 Multicalibration Post-Processing Algorithms and Hyperparameter Selection
In theory, standard calibration post-processing methods like Platt scaling (Platt et al., , 1999) or temperature scaling (Guo et al., , 2017) do not guarantee that predictions will be well-calibrated on protected subgroups. Therefore, in order to achieve multicalibration, Hébert-Johnson et al., (2018) propose an iterative boosting-style post-processing algorithm which we refer to as HKRR. The algorithm works by iteratively searching for and removing subgroup calibration violations until convergence. We detail the algorithm’s hyperparameters and the values we choose for them in Section D.1, and note that we perform a relatively wide parameter sweep.
The recent work of Haghtalab et al., (2023) also provide a family of alternative multicalibration algorithms with better theoretical sample complexity guarantees. This is motivated by the fact that HKRR is known to be theoretically sample inefficient (Gopalan et al., 2022b, ), and easily overfits (Detommaso et al., , 2024). For example, the number of samples required for generalization guarantees of HKRR is typically $O(\tfrac{1}{\alpha^{4}\lambda^{1.5}})$, where $\alpha$ determines the allowed multicalibration violation and $\lambda$ represents a suitable discretization width. For reasonably small values of $\alpha$ and $\lambda$, this can balloon the required number of samples to an unreasonable number. At a high level, each algorithm of Haghtalab et al., (2023) corresponds to a certain two-player game. Different algorithms in the family are a consequence of each player playing a different online learning algorithm. We detail the hyperparameters over which we search in Section D.2, but note here that we use the same code and predominantly the same parameters reported by the authors. We refer to any (post-processing) algorithm in this family as HJZ.
In addition to the multicalibration algorithms HJZ and HKRR, we test the usefulness of three standard calibration techniques in reducing multicalibration error: Platt scaling (Platt et al., , 1999), isotonic regression (Zadrozny and Elkan, , 2002), and temperature scaling Guo et al., (2017). The first two techniques are hyperparameter-free, and we use implementations given by Scikit-learn. We also use a parameter-free version of temperature scaling which we detail in Section D.3.
2.2 Subgroup Selection and Experimental Methodology
Multicalibration post-processing requires the selection of “groups” or subsets of the population of interest. As our investigation is primarily motivated by fairness desiderata, these subgroups determine what segments of the population the practitioner would like to “protect” or guarantee performance over. In most practical applications, these subgroups are constructed via features or conjunctions of features given as input for each data point. This way of constructing groups is standard: it is used by large production systems such as LinkedIn (Quiñonero Candela et al., , 2023), in the measurement of bias in ML (Tifrea et al., , 2024; Atwood et al., , 2024) and NLP systems (Baldini et al., , 2022; Li et al., , 2023), and in the auditing of large, deployed ML systems like Facebook (Ali et al., , 2019; Imana et al., , 2024). Furthermore, this approach still takes advantage of group intersectionality notions (Ovalle et al., , 2023), a cornerstone of multigroup fairness.
Remark 1
Using conjunctions of features or additional meta-data is not the only way to construct subgroups of the data. For any imperfect predictor $\hat{p}$, we can (nearly) always construct groups against which the predictor is not multicalibrated. For example, simply take the set of data points for which $\hat{p}$ predicts the incorrect label. Groups defined in this way may potentially be “as complex” as the underlying predictor $\hat{p}$. Nonetheless, it is not clear whether this is a meaningful set of groups to ask for multicalibration against (as it may require post-processing $\hat{p}$ to be a perfect predictor). To avoid such discussion, we intentionally determine groups by available features which we hypothesize practitioners may deem important or “sensitive” to the underlying prediction task.
We experiment across a variety of classification tasks: five tabular datasets (ACS Income, UCI Bank Marketing, UCI Credit Default, HMDA, MEPS), two language datasets (Civil Comments, Amazon Polarity), and two image datasets (CelebA, Camelyon17). For each dataset, we also define between 10 and 20 overlapping subgroups depending on available features or metadata. We detail and provide citations for each of our datasets and exact subgroup descriptions in Appendix C. In what follows, we give a high level overview of how we determined subgroups in our experiments.
For our tested tabular datasets, we determined groups by “sensitive” attributes—individual characteristics against which practitioners would not want to discriminate. In many cases, such attributes naturally include race, gender, and age, and vary with available information. For example, on ACS Income, we include groups such as “Multiracial,” “American Indian,” and “Seniors.” We also include some groups which are conjunctions of two attributes, for example “Black Women” or “White Women.” On Credit Default, for example, we include groups such as “Education = High School, Age > 40,” or “Married, Age > 60.” Similar groups are used throughout all tabular datasets, as well as CelebA.
On datasets where samples are not in correspondence with individuals—Camelyon17, Amazon Polarity, and Civil Comments—we define groups based on available information that can be viewed as “sensitive” with respect to the underlying task. In other words, we define groups such that an individual or institution using a predictor which is miscalibrated on this group may be seen as discriminating against the group. For example, a social media service should ideally not be underconfident when predicting the toxicity of posts mentioning a minority identity group; such predictions may allow hate speech to remain on the platform, or may provide differential engagement boosting based on the presence of racial identifiers in posts. Therefore, we include “Muslim,” “Black,” and various other phrases defining protected groups in the Civil Comments dataset. In the same vein, one would want a shopping platform to promote product listings fairly: a predictor which outputs mis-multicalibrated product ratings may boost listings for technology products proportionally to their true ratings but unfairly ignore the positive reviews of book listings. Thus, some groups from the Amazon Polarity dataset are defined as reviews containing the phrases “book,” “food,” “exercise,” or “tech.”
In Section 6.1, we further discuss group selection methodology and speculate about other ways of achieving multicalibration via group design.
Data Partitioning.
For consistency, we partition all datasets into three subsets: training, validation, and test. Test sets remain fixed across all experiments. We report accuracy and multicalibration metrics on the test set averaged over five random splits of train and validation sets for tabular data, and three splits for more complex data. Whenever a (multi)calibration post-processing algorithm is used, we run it using a holdout set of variable size from our training set, which we term the calibration set. We define the fraction of the training set used in (multi)calibration post-processing to be the calibration fraction. The exact calibration fractions over which we search appear in Section E.1 for tabular datasets and Section E.2 for image and language datasets. Note that multicalibration post-processing methods are far less sample efficient than standard post-hoc calibration methods. Therefore, the calibration fractions we test are broadly distributed between 5% and 100% of the training data (rather than using, say, a standard 10% of data for post-hoc calibration).
The calibration set is used solely in multicalibration post-processing, and is not used in training a model prior to the post-process. This procedure is motivated by a need to measure the importance of fresh samples in multicalibration post-processing. If a model is already multicalibrated on its entire training set $S$, we cannot re-use $S$ in HKRR or HJZ to improve the model, since the algorithms cannot improve on a predictor which is already perfectly multicalibrated on a particular dataset. This applies to models such as neural networks, which usually fit their training set to very low calibration error and high accuracy (Carrell et al., , 2022). For these models, we also anticipate that the multicalibration error on the training set will be low, and hence, the data from $S$ unusable for post-processing. [^3] Indeed, we test this more rigorously in Section 5, where we experiment with data reuse between model training and multicalibration post-processing. Therefore, the calibration fraction itself is an important hyperparameter we consider. Ideally, in order to maximize the resulting accuracy of the final model, we would utilize as much data as possible for model training, and minimize the amount of data required for multicalibration post-processing. However, due to the sample complexity of multicalibration algorithms, we will see that finding this specific point can be difficult (see Figure 3).
Compute. All experiments were performed on a collection four AWS G5 instances, each equipped with a NVIDIA 24GB A10 GPU. We used only the CPU for multicalibration and calibration post-processing, which was by far the most computationally intensive task. We estimate that all of our experiments cumulatively took 10 days of running time on these four instances.
3 Experiments on Tabular Datasets
We begin our investigation with tabular data. Although simpler than vision or language data, tabular data is an important and realistic setting which many algorithmic fairness practitioners encounter throughout the health, criminal justice, and finance sectors (Barda et al., , 2021; Obermeyer et al., , 2019; Barenstein, , 2019). As our base predictors in this setting, we consider multilayer perceptron NNs (MLPs), decision trees and random forests, SVMs, naive Bayes, and logistic regression. We defer dataset and group details to Appendix C, and model details to Section E.1. We note here that our datasets span from 10K examples (MEPS) to 200K (ACS Income), and we vary the size of the calibration set between 5% to 100% of the available training data. All of our results are computed with a mean and standard deviation over 5 train / validation splits. We instill the following insights from running multicalibration post-processing algorithms over 40,000 times on more than 1,000 separately trained models.
Observation 1: On tabular data, ML models which tend to be calibrated out of the box also tend to be multicalibrated without additional effort.
In Figure 1, we show the performance of every choice of multicalibration algorithm (corresponding to each choice of aforementioned hyperparameters) for MLPs on three datasets: MEPS, Credit Default, and ACS Income. We find that ERM performs nearly as well — in terms of worst group calibration error — as the best set of hyperparameters for HJZ and HKRR across our wide parameter sweep. This is seen broadly across all of our tabular datasets for models which one may expect to be calibrated in practice, such as logistic regression or random forests. [^4] We use the Scikit-learn random forest implementation, which predicts a probability corresponding to the fraction of positive points at the leaf. We include the complete plots of all multicalibration runs versus ERM in Section F.1.
We provide further evidence for Observation 1 by inspection of Figure 2. This table corresponds to the best choice of hyperparameters (according to maximum group-wise smECE on a validation dataset) of each method tested on the MEPS dataset. We find that HKRR and HJZ show no statistically significant improvements to max smECE for MLPs, random forests, and logistic regression. The gains offered by HKRR and HJZ in terms of worst group calibration error are also marginal (0 to 0.01) on the Bank Marketing, ACS Income, and Credit Default datasets (see Section F.2). On HMDA, however, multicalibration does seem to provide a noticeable improvement on the order of 0.03-0.07 for MLPs, random forests, and logistic regression (Figure 27). We believe this is because ERM achieves worse calibration error on HMDA, possibly due to the increased difficulty of the dataset.
Observation 2: HKRR or HJZ post-processing can help un-calibrated models like SVMs or naive Bayes achieve low group-wise maximum calibration error. Oftentimes, however, similar results can be achieved with traditional calibration methods like isotonic regression (Zadrozny and Elkan, , 2002).
Across our datasets, we find that SVMs, decision trees, and naive Bayes almost always have their max smECE error improve by 0.05 or more using multicalibration post-processing. We also point out the relatively strong performance of isotonic regression and other traditional calibration methods across datasets and models. For example, isotonic regression provides nearly all the improvements (up to 0.01 error) of the multicalibration algorithms when applied to naive Bayes in Figure 2. On the Credit Default dataset in Figure 28, isotonic regression is — when considering standard deviation — tied with the optimal multicalibration post-processing algorithms for SVM and naive Bayes. We have similar findings for the MEPS dataset and random forests trained on the HMDA dataset. Platt scaling and isotonic regression are desirable because they are parameter-free methods which work out of the box without tuning, are simple for practitioners to implement, and further do not require large parameter sweeps to find effective models.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.022 ± 0.006 0.106 ± 0.009 0.024 ± 0.002 0.086 ± 0.015 0.864 ± 0.001 MLP HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 MLP HJZ 0.019 ± 0.003 0.088 ± 0.011 0.021 ± 0.002 0.076 ± 0.018 0.864 ± 0.003 MLP Isotonic 0.02 ± 0.006 0.108 ± 0.021 0.02 ± 0.004 0.089 ± 0.021 0.864 ± 0.003 RandomForest ERM 0.019 ± 0.001 0.094 ± 0.006 0.021 ± 0.001 0.083 ± 0.004 0.863 ± 0.003 RandomForest HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 RandomForest HJZ 0.021 ± 0.004 0.106 ± 0.011 0.021 ± 0.003 0.101 ± 0.012 0.86 ± 0.003 RandomForest Isotonic 0.015 ± 0.002 0.089 ± 0.014 0.017 ± 0.001 0.084 ± 0.014 0.862 ± 0.002 SVM ERM 0.143 ± 0.002 0.376 ± 0.012 0.072 ± 0.001 0.186 ± 0.006 0.857 ± 0.002 SVM HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 SVM HJZ 0.031 ± 0.003 0.156 ± 0.021 0.027 ± 0.004 0.155 ± 0.02 0.828 ± 0.002 SVM Isotonic 0.048 ± 0.023 0.231 ± 0.085 0.048 ± 0.023 0.218 ± 0.069 0.847 ± 0.017 LogisticRegression ERM 0.022 ± 0.002 0.106 ± 0.008 0.022 ± 0.001 0.083 ± 0.003 0.866 ± 0.002 LogisticRegression HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 LogisticRegression HJZ 0.021 ± 0.003 0.114 ± 0.019 0.023 ± 0.001 0.09 ± 0.011 0.866 ± 0.003 LogisticRegression Isotonic 0.017 ± 0.003 0.109 ± 0.019 0.019 ± 0.003 0.097 ± 0.02 0.863 ± 0.002 DecisionTree ERM 0.067 ± 0.004 0.261 ± 0.028 0.047 ± 0.004 0.166 ± 0.012 0.85 ± 0.006 DecisionTree HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 DecisionTree HJZ 0.031 ± 0.003 0.156 ± 0.021 0.027 ± 0.004 0.155 ± 0.02 0.828 ± 0.002 DecisionTree Isotonic 0.014 ± 0.003 0.196 ± 0.026 0.015 ± 0.003 0.186 ± 0.027 0.838 ± 0.01 NaiveBayes ERM 0.277 ± 0.019 0.544 ± 0.02 0.164 ± 0.013 0.287 ± 0.011 0.714 ± 0.018 NaiveBayes HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 NaiveBayes HJZ 0.031 ± 0.003 0.156 ± 0.021 0.027 ± 0.004 0.155 ± 0.02 0.828 ± 0.002 NaiveBayes Isotonic 0.019 ± 0.005 0.128 ± 0.017 0.021 ± 0.005 0.122 ± 0.015 0.831 ± 0.006
Figure 2: Best performing HKRR and HJZ post-processing algorithm hyperparameters (selected based on validation max smECE) compared to ERM on the MEPS dataset. Calibrated models (MLP, random forest, logistic regression) need not be post-processed to achieve multicalibration. However, uncalibrated models (SVM, decision trees, naive Bayes) do benefit from multicalibration post-processing algorithms. Cells highlighted in blue show the importance of the choice of metric for selecting the best post-processing method for decision trees. Metric choice — worst group ECE vs. worst group smECE — can change which of ERM or HJZ is preferable.
Observation 3: A practitioner utilizing multicalibration post-processing can potentially face a trade-off between worst group calibration error and overall accuracy. This is most salient in high calibration fraction regimes (40-80%).
Due to the necessity of using hold-out data for running multicalibration post-processing, practitioners may have to choose between accuracy and worst group calibration error. For example, in Figure 3 we show that running multicalibration post-processing for MLPs on the HMDA dataset has a different optimal calibration fraction when considering accuracy and worst group calibration error as separate objectives. In fact, in this example, improving multicalibration error comes at a cost to accuracy of about 2%. Although small, this does indicate that the decision to use multicalibration post-processing here should be context-dependent. Additional examples of this tradeoff include decision trees or logistic regression on most datasets (Figures 31 and 33). We note, however, that these tradeoffs are more apparent in the higher calibration fraction regimes, where most of the training data is held out for multicalibration post-processing. This regime is potentially less relevant to practitioners, who usually reserve most of the available data for base model training. Plots for each dataset and model are in Section F.3.
Refer to captionRefer to captionRefer to caption
Figure 3: (Left/Middle): Hold-out calibration fraction vs. worst group calibration error (left) and accuracy (right) for MLPs on HMDA. The left plot shows that the best max group smECE is achieved at 60% of data used for base model training, and 40% of data for multicalibration post-processing. However, the right plot shows that the best accuracy is achieved by using 90% of data for base model training, and the remaining for multicalibration post-processing. This means that a practitioner may face a tradeoff between worst-group calibration error and accuracy. The impact of calibration fraction for each dataset is available in Section F.3. (Right): Gap between measured smECE and ECE on every group for every experiment. As sample size increases, the two metrics become very similar. However, some variability exists at lower sample sizes.
Observation 4: On small datasets, there can be variations between smECE and the standard binned ECE.
To illustrate an example, if a practitioner were selecting the best post-processing method for Decision Trees from Figure 2 based on ECE (see table cells highlighted in blue), HJZ may seem like a reasonable choice since it has a worst group ECE calibration error of 0.156. However, when using worst group smECE to measure performance, HJZ does not significantly improve upon ERM. This has an important consequence: if selecting the best model based on only the worst subgroup calibration error, the choice of calibration metric used will impact the choice of model.
In the rightmost plot of Figure 3, we also show each group’s sample size vs. the gap between measuring the group calibration error with smECE vs using ECE (over all datasets and groups). We find that as the group sample size increases, the gap between the metrics generally shrinks, and Observation 4 becomes less relevant. We note, however, that even on the ACS Income dataset with 200K examples, we find a significant difference of $0.1$ between measuring the overall calibration error of SVMs with ECE vs. smECE (Figure 25). More generally, to avoid issues stemming from ECE bin choice, we recommend that practitioners utilize the smECE calibration measurement tool [^5] https://github.com/apple/ml-calibration due to its theoretical guarantees and stability across our experiments.
Observation 5: When considering statistical significance, there is no clearly dominant algorithm between HKRR and HJZ on tabular data. However, HJZ is more robust towards the tested choice of hyperparameters. This may allow practitioners utilizing HJZ to find good solutions faster than using HKRR when post-processing simpler models such as naive Bayes or decision trees.
Over all tabular datasets and all base models (Section F.2), HJZ and HKRR had statistically distinguishable performance on 24 out of 30 cases. Among these 24 cases, HJZ performed better 7 times. Nonetheless, we observe that the HJZ family of algorithms is usually less sensitive to hyperparameter changes. In Figure 1 for example, the green points corresponding to hyperparameter choices for HJZ are tightly concentrated around ERM. We observe similar phenomena throughout additional model and dataset plots in Section F.1. Practitioners wishing to apply smaller hyperparameter searches over multicalibration algorithms may consider HJZ a suitable option, even if it gives slightly suboptimal worst group calibration error.
Additional Experiments. To understand how sensitive our observations are to our particular choice of group collection $G$, we also validate each of these observations with a new set of defined groups $G^{\prime}$, whose definitions are found in Section C.5. The full tabular results and plots for these new groups for each dataset is in Section G.1. Overall, the takeaway for most models (including MLPs) largely remains the same: it is difficult to find instances where multicalibration helps in a statistically significant way over ERM or some form of simple calibration (Observations 1 and 2). Further, where multicalibration does help, this help sometimes comes at a cost to accuracy (Observation 3).
4 Experiments on Language and Vision Datasets
In this section, we evaluate the ability of multicalibration post-processing to improve upon the multicalibration of vision transformers, DistilBERT, ResNets, and DenseNets on a collection of image and language tasks. Our goal is to understand if multicalibration post-processing can help in more complicated, large-model regimes within both the train-from-scratch and pre-trained paradigms.
As we move from smaller, tabular datasets to larger image and language datasets, we find that multicalibration algorithms may provide empirical improvements. Note here that in cases where we use a ResNet on language data, we train from scratch but use pretrained GloVe embeddings in the fashion of Duchene et al., (2023). In cases where we use a ResNet or DenseNet on image data, we also train from scratch. Whenever using a transformer, we finetune from pretrained weights (Sanh et al., , 2019; Dosovitskiy et al., , 2021). We defer further dataset, model, and group information to Appendix C and Section E.2.
Multicalibrating large models has an increased computational cost: For a single base predictor on tabular data, a full parameter sweep—training the multicalibration algorithm with every choice of hyperparameter—required 1-2 hours on a typical calibration set of 100K examples. With more complex base predictors (e.g. ResNets or language models) and larger datasets, this process takes significantly longer. Additionally, due to the increased computational cost of re-training a model in the image and language regimes, we only search over calibration fractions in $\{0.0,0.2,0.4\}$ and report our results averaged over three random train / validation splits. After running 1,700 multicalibration post-processing algorithms, we distill our findings into the following observations.
Observation 6: For image and language data, HKRR nearly always outperforms HJZ.
In all six of our experiments on image / language data (Section H.2), HKRR either matched or significantly beat the performance of HJZ. Note that we use the same parameter sweeps for HKRR and HJZ over image/language datasets that we used for the tabular datasets (see Appendix D), and leave open the possibility that HJZ may require a larger hyperparameter sweep to achieve good performance on more complex data.
Observation 7: On language and vision data, multicalibration post-processing can improve worst group calibration error relative to neural network ERM baselines by 50% or more. This stands in contrast to multicalibration post-processing for neural network ERM on tabular data, where we found no statistically significant improvements.
Over all language and vision datasets, HKRR improved worst group calibration error in 5 out of 6 cases. Among these 5, the least improvement we saw was HKRR decreasing the worst group smECE of ERM from 0.06 to 0.043 (DistilBERT on Civil Comments). The greatest improvement we saw was from 0.07 to 0.02 (ViT on Camelyon17) and 0.09 to 0.05 (ResNet-56 on Amazon Polarity). These examples all appear in Figure 4. A full collection of tables and plots can be found in Section H.2.
Observation 8: Binned ECE and smECE provide nearly identical estimates of calibration error.
Among nearly all of our experiments on larger datasets of more than 100k examples, we were not able to find any datasets where the metric used to measure calibration error would change the outcome of chosen model. This suggests that larger sample sizes close observable gaps between calibration measures (c.f. Observation 4).
Refer to captionRefer to captionRefer to caption
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ DistilBERT ERM 0.021 ± 0.001 0.065 ± 0.005 0.021 ± 0.001 0.06 ± 0.004 0.915 ± 0.001 DistilBERT HKRR 0.013 ± 0.0 0.047 ± 0.005 0.013 ± 0.0 0.043 ± 0.004 0.915 ± 0.001 DistilBERT HJZ 0.004 ± 0.001 0.043 ± 0.008 0.007 ± 0.001 0.043 ± 0.007 0.915 ± 0.001 DistilBERT Isotonic 0.002 ± 0.0 0.032 ± 0.006 0.005 ± 0.0 0.032 ± 0.006 0.916 ± 0.0 ResNet-56 ERM 0.039 ± 0.013 0.094 ± 0.009 0.039 ± 0.013 0.094 ± 0.009 0.867 ± 0.001 ResNet-56 HKRR 0.015 ± 0.001 0.059 ± 0.01 0.015 ± 0.001 0.047 ± 0.005 0.848 ± 0.004 ResNet-56 HJZ 0.013 ± 0.005 0.081 ± 0.012 0.014 ± 0.005 0.081 ± 0.012 0.863 ± 0.002 ResNet-56 Isotonic 0.005 ± 0.001 0.079 ± 0.009 0.007 ± 0.0 0.078 ± 0.008 0.863 ± 0.002
Figure 4: Top: Test accuracy vs. maximum group-wise calibration error (smECE) averaged over three train/validation splits for ViT and DenseNet on Camelyon17, and DistilBERT on CivilComments. Multicalibration post-processing has scope for improvement in each setting, and does so with nearly no loss in accuracy. Bottom: Impact of multicalibration post-processing algorithms for Civil Comments (DistilBERT) and Amazon Polarity (ResNet-56). Multicalibration post-processing and isotonic regression both offer improvements to worst group calibration error. Full results are available in Section H.1.
5 Simplifying Data Partitioning with Data Reuse
As discussed in Section 2.2, throughout our experiments on tabular, vision, and language data in Sections 3 and 4, we held out a portion of data from training solely for running multicalibration post-processing. This was motivated by two facts: (1) multicalibration may require fresh samples for theoretical statistical guarantees; and (2) if a model already has low worst group calibration error on a holdout set $S$, that set $S$ cannot be used for post-processing since there are no “group calibration violations” that either HJZ or HKRR can correct. Fact (1) holds for any base model (neural networks, decision trees, random forests, etc.), while (2) only holds for models which we believe may always achieve perfect training loss or calibration error, like neural networks.
We investigate these two facts experimentally by asking whether a practitioner may reuse a portion of the base model training data for multicalibration post-processing. Such data reuse could be very convenient to practitioners already saddled with hyperparameter optimizations. To test for this, we run ERM on all available training data, and run each post-processing method on this same set of data. Again searching over all post-processing hyperparameters (now except for calibration fraction), we provide test results corresponding to hyperparameters that achieved the best validataion max smECE. We provide full results for all tabular datasets in Section F.4.
Observation 9: Reusing model training data for multicalibration post-processing can sometimes be competitive with holding out data for post-processing. However, it can also come at a steep cost to worst group calibration error.
In some cases, reusing training data can marginally improve max smECE over the setup in which we do not reuse data; this is true, for example, for MLPs post-processed with HKRR on the left of Figure 5. Here, reusing training data improves upon using a calibration holdout by 0.025. In the vast majority of cases, however, reusing training data either (1) does not improve upon utilizing a holdout calibration set (for most models on ACSIncome or HMDA); or (2) significantly hurts (for MEPS, CreditDefault, or BankMarketing). For example, many base models are significantly hurt by data reuse for multicalibration post-processing on the CreditDefault dataset (see the right plot of Figure 5), having their post-processed performance drastically drop by 0.05-0.1 worst group smECE. These results demonstrate that practitioners utilizing multicalibration algorithms in practice may be required to optimize over the calibration fraction holdout size in order to achieve competitive empirical performance.
Figure 5: Impact of reusing all model training data for multicalibration post-processing on HMDA (Left) and CreditDefault (Right) as measured by worst group calibration error (max smECE). Results vary; for HMDA, post-processing with reused data essentially performs as well as post-processing by holding out data for all models except random forest postprocessed with HKRR. However, on CreditDefault, we find that data reuse can harm post-processing across the board. Plots for each dataset available in Section F.4.1.
6 Takeaways for Practitioners and Discussion
In this section, we first provide reasonable recommendations to practitioners wishing to apply multicalibration algorithms in practice. In Section 6.1, we also discuss additional details on the subgroup selection problem, which practitioners applying post-processing methods may find helpful.
We now present some takeaways that span all of our experiments. First, we believe that the latent multicalibration of ERM has been generally underestimated for many models. In particular, on tabular datasets, multicalibration post-processing cannot improve upon ERM for MLPs (see Observation 1). Furthermore, the improvement offered for more complex image and language data is generally less than 0.05 smECE when considering standard deviation.
This directly motivates our next takeaway: Current multicalibration post-processing algorithms—when applied to calibrated models like neural networks—are extremely sensitive towards choice of hyperparameters, since the potential “scope” of improvement is on the scale of 0.02 to 0.03 smECE. The optimal hyperparameter choice for each algorithm largely varies by dataset and base model, and it takes quite a bit of granular searching to find the best performing algorithm, or indeed, an algorithm which improves upon ERM at all. For example, the optimal HJZ algorithm used at least 15 different hyperparameter configurations across only our 30 tabular experiments (when considering calibration fraction as an additional parameter); HKRR has similar sensitivity issues. Further, many hyperparameter choices do not seem to improve upon the ERM base model—for example, see DenseNet-121 in Figure 4 or the full multicalibration post-processing plots in Section H.1—making a significant portion of the hyperparameter sweeps not useful to perform. Since training HJZ or HKRR on a holdout of 100K examples can take 1-2 hours, it can be several hours before a suitable choice of hyperparameters is found. This computational cost is exacerbated in the larger regimes where multicalibration may be most useful, which poses a major obstacle for practical applications of either HKRR or HJZ.
As a direct stopgap measure, we recommend running and evaluating traditional calibration methods. As we point out in Observation 2, post-processing algorithms like isotonic regression can achieve nearly the performance of multicalibration algorithms on tabular data. Isotonic regression also directly improves worst group calibration error over ERM in 4/6 of our experiments on larger models (see, e.g., Figures 4, H.2). Due to the fact that it is efficient and parameter free, we do not see a downside to running Isotonic regression (or any other calibration method) and testing if the maximum group-wise calibration error is beneath a desired threshold.
6.1 Subgroup Design Considerations
Subgroup selection and design is inhererently important to running multicalibration post-processing, given that choice of subgroup collection can drastically impact both the post-processed model and resulting calibration metrics. For practitioners, there are (at least) two important properties of groups to consider during the group selection phase: minimum group size and the richness of the group collection. The minimum group size $\gamma$ is a parameter which has implications for the overall sample complexity of multicalibration. In particular, it introduces a $1/\gamma$ factor into known sample complexity upper and lower bounds (Hébert-Johnson et al., , 2018; Shabat et al., , 2020). Note that $\gamma\in[0,1]$ is the size — as a fraction of the dataset under a distribution $\mathcal{D}$ — of the smallest group in the collection $G$. Therefore, there is a tradeoff between the size of the smallest group considered, and the number of samples needed for good multicalibration generalization.
In our experiments, we restricted to groups which were >0.5% of the entire dataset ($\gamma=0.005$). This was a reasonable “sweet spot” for us: Without enough samples from a particular group, known multicalibration algorithms are prone to overfitting the training set and not providing desirable generalization performance. On the other hand, if groups are not small enough, then the guarantees provided by running multicalibration algorithm may not be much better than standard calibration methods. Note that we consider collections of groups with sizes spanning from 0.5% all the way to 70-80% of the data (see Appendix C for detailed information). We deem this range reasonably sufficient to capture the varying sizes of groups that a practitioner may desire to protect in practice.
Group richness is also an important factor in group design. As discussed in Section 2.2, our results are only relevant to the setting where (1) we have well-defined groups that we seek to protect which are defined by simple features or conjunctions of features; and (2) we run multicalibration post-processing with those same groups. These two points naturally give rise to a potentially promising direction: is it practically feasible to multicalibrate with respect to a richer class of subgroups in order to obtain better empirical performance with respect to the simpler groups which one may actually care about? For example, one could test whether multicalibrating with respect to the group of all halfspaces would provide improved worst group calibration error over the class of feature-defined groups than one may actually care about final performance for. [^6] Surprisingly, it is possible to multicalibrate with respect to the (rich) class of all halfspaces with access to an agnostic halfspace learner (Hébert-Johnson et al., , 2018). In general, such agnostic learners are computationally hard to obtain in theory, but usually easy to construct in practice (via ERM). We leave such exploration to future work.
7 Experimental Limitations and Conclusion
One limitation of our results on language and vision datasets — as opposed to the tabular datasets — is that we do not optimize for the best choice of base model hyperparameters on each separate calibration split (e.g. changing learning rate or batch size between calibration fractions 0.2 and 0.4). In practice, we imagine that a practitioner would first fit the best possible ERM hyperparameters without using any holdout calibration data. Subsequently, they would then experiment with modifying the amount of hold out data to feed to multicalibration. We choose our hyperparameters for our base model to reflect this particular tuning strategy.
Another limitation of our results is that they are restricted to binary classification problems. While multicalibration algorithms do extend to multiclass problems, this extension comes at a severe cost of sample efficiency usually exponential in the number of labels (Zhao et al., , 2021). We show that — at least for tabular datasets — current multicalibration algorithms do not significantly improve upon a competitive and calibrated ERM baseline. If we were to further burden the multicalibration algorithm with the larger sample complexity of an additional label, we do not expect that their performance will improve. Nonetheless, we plan to investigate the multiclass setting in future work, and believe that those findings will be consistent with the results present in this paper.
Finally, we do not offer much explanation of why we see differing performance of the two algorithms HJZ and HKRR; we offer some discussion of this in Section A.3, but more is warranted in future work.
We believe that our work illuminates many avenues towards improving the viability of multicalibration algorithms in practice. For example, developing parameter free multicalibration methods (akin to what smECE accomplishes for calibration metrics) is an important direction with direct impacts on the practice of fair machine learning. Similarly, post-processing techniques with better empirical sample complexity could significantly help the practice of multicalibration.
Acknowledgements.
SD was supported by the Department of Defense through the National Defense Science & Engineering Graduate (NDSEG) Fellowship Program. This work was also supported by NSF CAREER Award CCF-2239265 and an Amazon Research Award. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of sponsors such as Amazon or NSF. The authors would like to thank Bhavya Vasudeva for discussions that were helpful in the design of early experiments, and Eric Zhao for help in utilizing the HJZ algorithms. The authors also sincerely thank the anonymous Neurips reviewers for providing detailed feedback and discussion which greatly improved and clarified parts of this work.
Appendix A Additional Related Work
Limitations of Calibration. A collection of works characterize the limitations of calibration as a property of predictors. Of particular note is Perez-Lebel et al., (2023), who remark that calibration is often misunderstood in the literature, as it does not guarantee that output probabilities are close to the ground truth probability distribution. We make no such claim about calibration, and only justify its use in order to ensure model predictions are meaningful. Yuksekgonul et al., (2024) draw connections between calibration and atypicality of certain examples, improving group wise-performance of NNs without subgroup annotations via what they term “atypicality-aware recalibration.”
Applications of Multicalibration. Beyond classification, Globus-Harris et al., (2023) introduce algorithms that post-process multicalibrated regression functions to satisfy a variety of fairness constraints, and present experiments using such algorithms on logistic regression and gradient-boosted decision trees. Benz and Rodriguez, (2023) study predictive confidence in the setting of AI-assisted decision making; they show the existence of distributions under which a “rational” decision maker is unlikely to find an optimal policy using calibrated confidence values, and prove that multicalibration with respect to the decision maker’s preliminary confidence values is often sufficient for aversion of such issues. Zhang et al., (2024) also utilize a generalization of multicalibration to tackle interesting problems like de-biased text generation and false negative rate control.
Calibration of NNs. Literature on the calibration of neural networks (NNs) is very rich; see for example a treatment by Wang, (2023). Most importantly, Minderer et al., (2021) have run comprehensive, large-scale experiments detailing the degree to which modern NNs are calibrated. Their results show that current state-of-the-art models appear to be nearly perfectly calibrated, and appear to remain so even in the presence of distribution shift. This is in somewhat striking contrast to the earlier results of Guo et al., (2017), which demonstrated the best models at time of their publication to be quite miscalibrated, and highlighted the need to further investigate calibration measures. Our focus in this work is instead on evaluating multicalibration of predictors on datasets over which we can naturally define a collection of protected subgroups.
Carrell et al., (2022) examine a connection between generalization and calibration generalization, the difference in calibration error on train and test sets. In particular, they claim DNNs to be well-calibrated on their training sets and the accuracy generalization gap to upper bound the calibration generalization gap. Such observations imply NNs that generalize well to be well-calibrated.
Trainable Calibration Measures. Laplace Kernel calibration measures have been shown to be effective in enforcing confidence calibration for neural networks when used in training. In particular, using MMCE (Maximum Mean Calibration Error) as a regularizer in tandem with cross-entropy loss yields high accuracy predictions while moderately improving calibration by taming overconfident predictions (Kumar et al., , 2018). Additionally, this metric is efficiently computable in quadratic time.
Subgroup Robustness. Several works in subgroup robustness literature examine the performance of NNs by worst-group-accuracy, particularly in cases where NNs tend to rely on spurious correlations. Recent works by Kirichenko et al., (2023); LaBonte et al., (2023) propose last-layer fine-tuning as a simple and computationally inexpensive way to do exactly that. Indeed, Mao et al., (2023) extend the method to address fairness concerns, appending a fairness constraint to the training objective during fine-tuning. Such works examine only one sensitive attribute at a time, and often consider the disjoint groups produced by unique values of this attribute in conjunction with label. Rosenfeld and Garg, (2023) show connections between robustness and distribution shift for neural networks via unlabeled test data.
In general, we view multi-group robustness as a notion of robustness which, like multicalibration, aims to “respect” multiple groups simultaneously. Perhaps the closest connection between multigroup robustness and multicalibration is: they intuitively may share similar mechanisms (as the reviewer noted). That is, the mechanistic question in both cases is to understand which groups can be easily computed from the predictor’s "features" – ie, which groups are easy for the predictor to distinguish.
A.1 Further Discussion of Błasiok et al., (2024)
The connection to our work is subtle. In Błasiok et al., (2024), the authors consider performing ERM over some $\mathcal{C}^{\prime}\supseteq\mathcal{C}$ so that the resulting model is multicalibrated with respect to $\mathcal{C}$. In their case, $\mathcal{C}^{\prime}$ is a family of neural networks (NNs), and $\mathcal{C}$ is some family of smaller NNs. In many of our experiments, we indeed perform (approximate) ERM over some family of NNs, and one might expect this to result in multicalibration with respect to our finite collection of groups $G$, which are easily computable by some class of smaller NNs. This is because for NNs, it is possible to represent arbitrarily-complex groups by simply taking "large enough" networks, without making design choices specific to the group structure.
In our tabular experiments, we restrict our experiments to small multi-layer perceptron networks of 3-4 layers, whose smaller subnetworks may not be learning complex functions capturing groups of interest (see Appendix E for model descriptions). Nonetheless, we find that these models possess latent multicalibration properties, as discussed further in Observation 1 in Section 3. Our vision and language experiments show that post-processing does have positive impact, suggesting that the models sub-networks are not sufficiently capturing the groups of interest (see Section 4).
A.2 Equivalence of Early Stopping and Hyperparameter Sweeps
Detommaso et al., (2024) utilize two early-stopping criterion. The first is to halt further multicalibration iterations once the group conditioned on the bin set becomes “too small.” We also utilize this technique, which is inherent in the algorithm of HKRR. Detommaso et al., (2024) also use a holdout validation set to early-stop a variant of HKRR when the mean-squared error (MSE) on the hold-out set fails to decrease. We instead vary the permitted violation parameter $\alpha$ of HKRR, and select the best parameter with a holdout validation set. $\alpha$ controls the permitted calibration error conditioned on a group and particular bin (of width $\lambda=0.1$ in our work). We believe that early stopping for a validation metric should have similar performance to running the full HKRR with a variety of $\alpha$ levels, and selecting based on validation performance. Nonetheless, we note that in our experiments, we select based on validation smoothECE multicalibration error, while Detommaso et al., (2024) select on validation MSE. This could potentially lead to performance differences.
A.3 Discussion on Dynamics and Performance of HJZ and HKRR
The dynamics of both of the algorithms HJZ and HKRR are complex. In particular, the performance of the algorithms depend on at least the following parameters:
  1. 1.
    Distribution of initial predictions output by the models (i.e. input to post-processing algorithm);
  2. 2.
    Choice of hyperparameters for HKRR and HJZ;
  3. 3.
    "Complexity” or “expressiveness” of the groups; and
  4. 4.
    Number of available samples, and whether the samples are re-used from training or not.
In our work, we focus mainly on (2) and (4). We discuss (3) to an extent in Section 6.1, but teasing apart exactly how (1) and (3) contribute to the performance of multicalibration post-processign algorithms is certainly an interesting avenue for future work. We believe that part of the reason for the superiority of HKRR in language/vision data may be explained within the lens of (2) and (4). That is, we may have found better hyperparameters for HKRR with a wider search, and the sample complexity may be better in practice than the game-theoretic approach offered by HJZ.
Due to computational constraints and the added dimension of choosing how much data to save for calibration, we search a large — but not all-encompassing — collection of hyperparameters for each of the multicalibration algorithms tested. With regards to dataset size, 3 of the 4 vision/language datasets are noticeably larger than the tabular datasets (by at least 100k samples). It is possible that HKRR generally performs better on such dataset sizes, or that the optimal hyperparameters for HJZ change significantly in this larger-sample regime.
Understanding why HJZ may be more stable to hyperparameter choices (see Observation 5 in Section 3) is a more challenging question to answer, since it likely has to do with internal game dynamics in the learning algorithm. In particular, by choosing various online learning algorithms, HJZ implements a family of multicalibration post-processing methods. We test all algorithms from this family with varying parameters. It is possible that the family of algorithms itself somehow has a shift in stability as we scale to a large data regime (4). However, as analyzing even a singular algorithm (e.g. HKRR) is challenging, we are not sure that speculating about the stability of a family of algorithms is currently possible, and hence, leave this to future work.
Appendix B Broader Impacts
Our work performs a comprehensive empirical evaluation of multicalibration post-processing, which could help practitioners apply these notions more effectively in practice. We note however, that fairness can be subtle, and multicalibration by itself may not be enough to ensure fairness. By now it is well understood that there are tradeoffs between different notions of fairness, and the right definitions to be deployed are context-dependent and depend on societal norms. Therefore, our results on the latent multicalibration of neural networks should not be construed as implying that these models are already fair, and care should be taken before deploying any ML model in applications with consequential societal outcomes.
Appendix C Dataset and Subgroup Descriptions
Here, we detail the datasets and group information used in all experiments.
C.1 Tabular Datasets
The ACS Income dataset, introduced by Ding et al., (2021), is a superset of the UCI Adult [^7] For comparison with prior work, we include the original UCI Adult dataset in our benchmark repository. dataset (Becker and Kohavi, , 1996) derived from additional US Census data. We use the folktables package introduced alongside the work. In particular, we consider the task of predicting whether an American adult living in California receives income greater than \$50,000 in the year 2018. Features include race, gender, age, and occupation. For this task, the dataset furnishes just under 200,000 samples.
The UCI Bank Marketing dataset documents 45,000 phone calls made by a Portuguese banking institution over the course of several marketing campaigns (Moro et al., , 2012). We consider the task of predicting whether, on a given call, the client will subscribe a term deposit, given features characterizing the housing, occupation, education, and age of the client.
The UCI Default of Credit Card Clients dataset (termed “Credit Default” in our experiments) documents the partial credit histories of 30,000 Taiwanese individuals (Yeh, , 2016). We consider the task of predicting whether an individual will default on credit card debt, given payment history and additional identity attributes.
The HMDA (Home Mortgage Disclosure Act) dataset documents the US mortgage applications, identity attributes of associated applicants, and the outcome of these applications (Federal Financial Institutions Examination Council, , 2017). We use a 114,000-sample variant of this dataset given by Cooper et al., (2023), and consider the task of predicting whether a 2017 application in the state of Texas was accepted.
The MEPS (Medical Expenditure Panel Survey) dataset comes from the US Department of Health and Human Services and documents healthcare utilization of US households. We use a 11,000-sample variant of the dataset, originally studied in Sharma et al., (2021), and consider the task of predicting whether a household makes at least 10 medical visits, given socioeconomic and geographic information of household applicants.
C.2 Image Datasets
The CelebA dataset, introduced by Liu et al., (2015), consists of 200,000 cropped and aligned images of celebrity faces. We consider the task of predicting hair color, a task known to be difficult for certain label-dependent subgroups due to the existence of spurious correlations (Sagawa et al., , 2019). Metadata documents certain characteristics of the individuals in the images such as gender, face shape, hair style, and the presence of fashion accessories.
The Camelyon17 dataset, introduced by Bándi et al., (2019), consists of histopathological images of human lymph node tissue. We use a patch-based variant of this dataset, introduced by Koh et al., (2021), which consists of 450,000 $96$x96969696 images. Unlike Koh et al., (2021), we shuffle the predetermined training and test splits. We consider the task of predicting whether a given image contains tumorous tissue. Metadata documents the hospital from which a given patch originates, and the original slide from which the patch is drawn.
C.3 Language Datasets
The CivilComments dataset, introduced by Borkan et al., (2019), contains 450,000 online comments annotated for toxicity and identity mentions by crowdsourcing and majority vote. We use the WILDS variant of this dataset, provided by Koh et al., (2021), though we shuffle the predetermined training and test splits, and consider the task of prediction whether a given comment is labeled toxic.
The Amazon Polarity dataset, also introduced by Zhang et al., (2015) and a subset of the Amazon Reviews dataset, provides the text content of 4,000,000 Amazon reviews. A review receives the label 1 when associated with a rating greater than or equal to 4 stars, and a label of 0 when associated with a rating of less than or equal to $2$ stars. As with Yelp Polarity, this dataset comes with no metadata, so we define groups based on the presence of meaningful words. We use a randomly-drawn, 400,000 sample subset across all experiments.
C.4 Groups for Tabular Datasets
Here we present the subgroups considered for each dataset in our experiments. In all cases, we only consider subgroups composing at least a 0.005-fraction of the underlying dataset.
Note that the ‘Dataset’ row in each table does not correspond to a group used in multicalibration post-processing, nor are aggregate metrics used to compute worst-group metrics such as max smECE. We include this row for convenience.
{adjustwidth}
-1in-1in group name n samples fraction y mean Black Adults 8508 0.0435 0.3461 Black Females 4353 0.0222 0.3193 Women 92354 0.4720 0.3491 Never Married 68408 0.3496 0.2344 American Indian 1294 0.0066 0.2836 Seniors 14476 0.0740 0.5410 White Women 55856 0.2855 0.3729 Multiracial 8206 0.0419 0.3572 Asian 32709 0.1672 0.4805 Dataset 195665 1.0000 0.4106
Figure 6: ACS Income groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Job $=$ Management 9458 0.2092 0.1376 Job $=$ Technician 7597 0.1680 0.1106 Job $=$ Entrepreneur 1487 0.0329 0.0827 Job $=$ Blue-Collar 9732 0.2153 0.0727 Job $=$ Retired 2264 0.0501 0.2279 Marital $=$ Married 27214 0.6019 0.1012 Marital $=$ Single 12790 0.2829 0.1495 Education $=$ Primary 6851 0.1515 0.0863 Education $=$ Secondary 23202 0.5132 0.1056 Education $=$ Tertiary 13301 0.2942 0.1501 Housing $=$ Yes 25130 0.5558 0.0770 Housing $=$ No 20081 0.4442 0.1670 Age $<$ 30 3050 0.0675 0.1951 30 $\leq$ Age $<$ 40 17359 0.3840 0.1129 Age $\geq$ 50 12185 0.2695 0.1287 Dataset 45211 1.0000 0.1170
Figure 7: Bank Marketing groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Male, Age $<$ 30 3281 0.1094 0.2405 Single 15964 0.5321 0.2093 Single, Age $>$ 30 6888 0.2296 0.1992 Female 18112 0.6037 0.2078 Married, Age $<$ 30 1482 0.0494 0.2611 Married, Age $>$ 60 225 0.0075 0.2667 Education $=$ High School 4917 0.1639 0.2516 Education $=$ High School, Married 2861 0.0954 0.2635 Education $=$ High School, Age $>$ 40 2456 0.0819 0.2577 Education $=$ University, Age $<$ 25 1610 0.0537 0.2795 Female, Education $=$ University 8656 0.2885 0.2220 Education $=$ Graduate School 10585 0.3528 0.1923 Female, Education $=$ Graduate School 6231 0.2077 0.1814 Dataset 30000 1.0000 0.2212
Figure 8: Credit Default groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Applicant Ethnicity: Hispanic or Latino 26416 0.2313 0.6806 Applicant Ethnicity: Not Hispanic or Latino 73527 0.6439 0.7940 Applicant Ethnicity: Not provided 14128 0.1237 0.6704 Applicant Sex: Female 32143 0.2815 0.7319 Applicant Sex: Male 72635 0.6361 0.7713 Co-Applicant Sex: Female 35164 0.3080 0.8029 Co-Applicant Sex: Male 10336 0.0905 0.7767 Applicant Race: Black 9044 0.0792 0.6703 Applicant Race: Asian 8086 0.0708 0.8097 Applicant Race: Native American or Alaskan 1019 0.0089 0.5927 Co-Applicant Race: Black 2760 0.0242 0.7120 Co-Applicant Race: Asian 3339 0.0292 0.8194 Dataset 114185 1.0000 0.7524
Figure 9: HMDA groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Age 0-18 3308 0.2986 0.0605 Age 19-34 2468 0.2228 0.1021 Age 35-50 2186 0.1973 0.1404 Age 51-64 1813 0.1636 0.2670 Age 65-79 977 0.0882 0.4637 Not White 7121 0.6427 0.1227 Northeast 1553 0.1402 0.2260 Midwest 2020 0.1823 0.2040 South 4325 0.3904 0.1487 West 3181 0.2871 0.1481 Poverty Category 1 2435 0.2198 0.1577 Poverty Category 2 704 0.0635 0.1378 Poverty Category 3 1941 0.1752 0.1484 Poverty Category 4 3100 0.2798 0.1519 Dataset 11079 1.0000 0.1694
Figure 10: MEPS groups.
C.5 Alternate Groups for Tabular Datasets
To validate our observations on tabular data, we repeated all experiments on each tabular dataset with an alternate collection of groups. Like the original groups, these groups are defined by a feature or conjunction of two features. We provide the alternate group definitions here, and present results on these groups in Appendix G.
{adjustwidth}
-1in-1in group name n samples fraction y mean Associates Degree Male 7331 0.0375 0.4957 Associates Degree Female 8372 0.0428 0.3186 Divorced Female 10652 0.0544 0.4415 Under Part Time 16525 0.0845 0.1025 Part Time 55269 0.2825 0.1408 Full Time 135989 0.6950 0.5214 Over Full Time 11471 0.0586 0.6441 Not White 74659 0.3816 0.3574 Government Employee 29121 0.1488 0.5337 Private Employee 166544 0.8512 0.3890 Under 21 10166 0.0520 0.0106 Middle Aged 81582 0.4169 0.5064 Dataset 195665 1.0000 0.4106
Figure 11: ACS Income alternate groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Job $=$ Management, Age $<$ 50 7091 0.1568 0.1393 Job $=$ Technician, Age $<$ 30 436 0.0096 0.1858 Job $=$ Blue-Collar, Age $>$ 50 1075 0.0238 0.0679 Married, Education $=$ Primary 5246 0.1160 0.0755 Single, Education $=$ Tertiary 4792 0.1060 0.1836 Housing $=$ Yes, Age $<$ 30 1621 0.0359 0.0993 Housing $=$ No, Age $<$ 30 1429 0.0316 0.3037 Under 21 305 0.0067 0.3115 Middle Age 23841 0.5273 0.0977 Senior Age 961 0.0213 0.4225 Dataset 45211 1.0000 0.1170
Figure 12: Bank Marketing alternate groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Single, Male 6553 0.2184 0.2266 Single, Female 15964 0.5321 0.2093 Young Adult 9618 0.3206 0.2284 Middle Aged 8872 0.2957 0.2356 Education $=$ High School, Female 2927 0.0976 0.2364 Education $=$ University, Female 8656 0.2885 0.2220 Education $=$ High School, Single 1909 0.0636 0.2368 Education $=$ High School, Married 2861 0.0954 0.2635 Education $=$ University, Single 7020 0.2340 0.2306 Education $=$ University, Married 6842 0.2281 0.2435 Education $=$ Graduate, Single 6809 0.2270 0.1842 Dataset 30000 1.0000 0.2212
Figure 13: Credit Default alternate groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Loan Type 1 87857 0.7694 0.7327 Loan Type 2 17587 0.1540 0.8125 Loan Type 3 8047 0.0705 0.8311 HUD Median Family Income $>$ 50k 107450 0.9410 0.7583 HUD Median Family Income $\leq$ 50k 6735 0.0590 0.6578 Has Co-Applicant 42535 0.3725 0.8079 Agency $=$ OCC 5966 0.0522 0.8235 Agency $=$ FRS 3879 0.0340 0.8729 Agency $=$ FDIC 6951 0.0609 0.8708 Agency $=$ NCUA 10626 0.0931 0.6882 Agency $=$ HUD 64915 0.5685 0.7562 Agency $=$ CFPB 21848 0.1913 0.6939 Loan Type $=$ 1 to 4 Family 87857 0.7694 0.7327 Loan Type $=$ Manufactured Housing 17587 0.1540 0.8125 Loan Type $=$ Multi-Family 8047 0.0705 0.8311 Dataset 114185 1.0000 0.7524
Figure 14: HMDA alternate groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Under 21 3772 0.3405 0.0607 Middle Age 2874 0.2594 0.1990 Senior Age 1304 0.1177 0.4862 Sex $=$ 1 5281 0.4767 0.1274 Sex $=$ 2 5798 0.5233 0.2077 White 3958 0.3573 0.2534 Active Duty Group 2 6454 0.5825 0.1432 Marriage Group 1 3645 0.3290 0.2222 Marriage Group 2 450 0.0406 0.4867 Pregnancy Group 1 124 0.0112 0.4274 Pregnancy Group 2 2167 0.1956 0.1398 Insurance Group 1 5926 0.5349 0.1790 Insurance Group 2 3890 0.3511 0.1979 Dataset 11079 1.0000 0.1694
Figure 15: MEPS alternate groups.
C.6 Groups for Image Datasets
{adjustwidth}
-1in-1in group name n samples fraction y mean Male 84434 0.4168 0.0207 Female 118165 0.5832 0.2389 Arched Eyebrows 54090 0.2670 0.2227 Bangs 30709 0.1516 0.2310 Big Lips 48785 0.2408 0.1629 Chubby 11663 0.0576 0.0189 Double Chin 9459 0.0467 0.0248 Eyeglasses 13193 0.0651 0.0392 High Cheekbones 92189 0.4550 0.1949 Mouth Slightly Open 97942 0.4834 0.1737 Oval Face 57567 0.2841 0.1761 Pale Skin 8701 0.0429 0.2455 Receding Hairline 16163 0.0798 0.0630 Smiling 97669 0.4821 0.1812 Straight Hair 42222 0.2084 0.1518 Wavy Hair 64744 0.3196 0.2145 Wearing Hat 9818 0.0485 0.0168 Young 156734 0.7736 0.1581 Dataset 202599 1.0000 0.1480
Figure 16: CelebA groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean Hospital $=$ 0 59436 0.1304 0.5000 Hospital $=$ 1 34904 0.0766 0.5000 Hospital $=$ 2 85054 0.1865 0.5000 Hospital $=$ 3 129838 0.2848 0.5000 Hospital $=$ 4 146722 0.3218 0.5000 Slide $=$ 0 4316 0.0095 0.0083 Slide $=$ 4 7294 0.0160 0.6697 Slide $=$ 8 13455 0.0295 0.6469 Slide $=$ 16 4971 0.0109 0.8236 Slide $=$ 20 3810 0.0084 0.0071 Slide $=$ 24 7727 0.0169 0.0238 Slide $=$ 28 31878 0.0699 0.8469 Slide $=$ 32 8831 0.0194 0.2466 Slide $=$ 36 10661 0.0234 0.0015 Slide $=$ 40 7395 0.0162 0.0170 Slide $=$ 44 7958 0.0175 0.0030 Slide $=$ 48 61110 0.1340 0.9273 Dataset 455954 1.0000 0.5000
Figure 17: Camelyon17 groups.
C.7 Groups for Language Datasets
{adjustwidth}
-1in-1in group name n samples fraction y mean Male 20880 0.0466 0.1488 Female 33113 0.0739 0.1399 LGBTQ 14303 0.0319 0.2684 Not LGBTQ 433695 0.9681 0.1083 Christian 18961 0.0423 0.1103 Not Christian 380222 0.8487 0.1177 Muslim 13939 0.0311 0.2429 Not Muslim 418737 0.9347 0.1065 Other Religions 11030 0.0246 0.1528 Black 8448 0.0189 0.3638 Not Black 426444 0.9519 0.1049 White 14339 0.0320 0.3068 Not White 415090 0.9265 0.1016 Dataset 447998 1.0000 0.1134
Figure 18: Civil Comments groups.
{adjustwidth}
-1in-1in group name n samples fraction y mean expensive 5834 0.0146 0.4434 cheap 10928 0.0273 0.2753 food 3868 0.0097 0.5476 health 2381 0.0060 0.6237 music 26463 0.0662 0.6192 book 108100 0.2703 0.5289 movie 36191 0.0905 0.4731 tech 8515 0.0213 0.4547 exercise 2262 0.0057 0.5535 garbage 3248 0.0081 0.0702 terrible 6138 0.0153 0.0893 incredible 2532 0.0063 0.7986 love 53005 0.1325 0.7212 again 26106 0.0653 0.4454 star 42632 0.1066 0.4495 Dataset 399980 1.0000 0.5009
Figure 19: Amazon Polarity groups.
C.8 Dataset Usage and Licensing
  • ACSIncome: While Folktables provides API for downloading ACS data, usage of this data is governed by the terms of use provided by the Census Bureau. For more information, see https://www.census.gov/data/developers/about/terms-of-service.html.
  • BankMarketing: Creative Commons Attribution 4.0 International (CC BY 4.0)
  • CreditDefault: Creative Commons Attribution 4.0 International (CC BY 4.0)
  • HMDA: The variant we use is available for download on https://github.com/pasta41/hmda?tab=readme-ov-file under an MIT license.
  • MEPS: The variant we use is available for download on https://github.com/alangee/FaiR-N/tree/master under an Apache 2.0 license.
  • Civil Comments: This dataset is in the public domain and distributed under CC0.
  • Amazon Polarity: We were not able to find a license for this dataset. It is a downstream variant of a dataset generated with content from Internet Archive [^8] http://archive.org/details/asin_listing/.
  • CelebA: The creators of this dataset do not provide a license, though they encourage its use for non-commercial research purposes only.
  • Camelyon17: This dataset is in the public domain and distributed under CC0.
Appendix D Hyperparameters for Multicalibration and Calibration Algorithms
Here, we detail the hyperparameters with which equip algorithms from Haghtalab et al., (2023) and Hébert-Johnson et al., (2018), as well as the hyperparameters used in standard calibration methods.
D.1 Hébert-Johnson et al., (2018) Algorithm
These authors do not report empirical performance of their algorithm. For consistency with calibration measures, we fix a bin width of $\lambda=0.1$ in all experiments. The algorithm depends on one other hyperparameter $\alpha$, which is some constant factor of the acceptable difference between mean prediction and mean label within each “category,” a subgroup $g\subseteq\mathcal{X}$ restricted to the preimage of a particular bin $b\subseteq[0,1]$. Modulo the randomness induced by a statistical query oracle, this algorithm converges when violations within each category are sufficiently small. As originally proposed, we skip categories $g\cap f^{-1}(b)$ in iterations where $|g\cap f^{-1}(b)|\leq\lambda\alpha|g|$. For each dataset and each base predictor, we sweep over $\alpha\in\{0.1,0.05,0.025,0.0125\}$.
D.2 Haghtalab et al., (2023) Algorithms
These authors present an empirical examination in conjunction with theoretical results, though our use of the algorithms differs significantly. Namely, instead of initializing predictions uniformly, we initialize with the predictions of some base predictor. The authors also train their multicalibration algorithms with substantially larger collections of subgroups, in some cases defining group by all unique values of individual features. Instead, we only consider a collection of at most 20 “protected” groups.
[…]
In all experiments, we consider the same multicalibration algorithms but restrict to 30 iterations and a smaller collection of decay rates. In particular, for each dataset and base predictor, and for each of the six algorithms, we sweep over decay rates $\eta\in\{0.9,0.95\}$ for the learner and $\eta\in\{0.9,0.95,0.98\}$ for the adversary. We justify these restrictions by noting that (1) our base predictors already achieve nontrivial accuracy on each task and (2) that this sweep covers a large portion of the optimal hyperparameters found in Haghtalab et al., (2023). For consistency with our hyperparameters for HKRR and chosen calibration measures, we fix $\lambda=0.1$ on all datasets.
D.3 Temperature Implementation
Temperature scaling can be made hyperparameter-free by choosing a divisor $T^{*}$ which minimizes the cross-entropy loss on a held-out calibration split. One can also fix $T$ to some positive real number. For each dataset and base predictor, we examine both methods, scaling logits by $1/T$ for all $T\in\{0.2\cdot k:k\in[20]\}$, as well as by $1/T^{*}$ with $T^{*}$ obtained via the Pytorch implementation of L-BFGS. We report only the best temperature scaling method on a hold-out validation set.
Appendix E Models and Training
Here, we describe the models used as base predictors and their hyperparameters (or the procedure for obtaining these hyperparameters) in all experiments. Across all datasets, we use a train-validation-test split of $(0.6:0.2:0.2)$, fixing the test set and determining train/validation sets via random seed. In all cases of a base-predictor hyperparameter search, we use validation accuracy to select hyperparameters.
E.1 Models Used on Tabular Data
On all tabular datasets, we examine five standard prediction models from supervised learning: Decision Tree, Random Forest, Logistic Regression, SVM, and Naive Bayes, using a Scikit-learn implementation in all cases. We also examine MLPs of varying architecture. For each dataset and prediciton model, we examine all calibration fractions $\operatorname{CF}\in\{0,0.01,0.05,0.1,0.2,0.4,0.6,0.8,1.0\}$. When $\operatorname{CF}=1.0$, we take the base predictor to output $1/2$ for all samples. During both hyperparameter search and test-set evaluation we average all metrics over five random splits of the training and validation data.
E.1.1 Standard Supervised Learning Models
For Decision Tree, we vary maximum depth over $\{\texttt{None},10,20,50\}$ and the minimum number of samples required to split an internal node over $\{2,5,10\}$. For Random Forest, we vary these same hyperparameters and fix the number of estimators at $100$. For Logistic Regression and SVM, we vary regularization strength; for Logistic Regression we let the inverse regularization strength $C\in\{0.4,1,2,4\}$ and for SVM we let the regularization strength $\alpha\in\{0.00001,0.0001,0.001,0.01\}$. Naive Bayes is hyperparameter-free.
While Decision Tree, Random Forest, Logistic Regression, and Naive Bayes are naturally probabilistic, SVM is not. While Scikit-learn provides a probabilistic prediction method with the predict_proba() function, this is implemented via Platt scaling of the SVM scores. For this reason, we treat the SVM’s standard output labels as probabilities. For efficient on larger datasets, we also use the SGDClassifier implementation of SVM.
E.1.2 MLPs
To reduce computation during MLP hyperparameter search, for each dataset we constructed a smaller set of hyperparameters over which to sweep, based on what yielded the best performance in preliminary experiments. In what follows, we let a sequence $(\ell_{i})_{i=1}^{N}$ denote the ordered layer widths for a particular MLP with $N$ hidden layers. When we substitute some $\ell_{i}$ with $\operatorname{BN}$, this indicates the presence of a batch-normalization layer. In all experiments with MLPs on tabular datasets, we use the Adam optimizer.
On ACS Income, we train for $50$ epochs. We search over hidden-layer widths: $(128,\operatorname{BN},128)$, $(128,256,128)$, and $(128,\operatorname{BN},256,\operatorname{BN},128)$. We vary batch size over $\{32,64,128\}$ and learning rate over $\{0.01,0.001,0.0001,0.00001\}$.
On Bank Marketing, we train for $50$ epochs. We search over hidden-layer widths of $(100)$, $(128,\operatorname{BN},128)$, $(128,256,128)$, and $(128,\operatorname{BN},256,\operatorname{BN},128)$. We vary batch size over $\{64,128,256,512\}$ and learning rate over $\{0.001,0.0001,0.00001\}$. We also include a learning rate schedule under which our learning rate is $0.00005$ for the first five epochs and $0.00001$ for the remaining.
On Credit Default, we train for $5$ epochs. We search over hidden-layer widths of $(100)$, $(128,256,128)$, and $(128,\operatorname{BN},256,\operatorname{BN},128)$. We vary batch size over $\{16,32,64,128\}$ and learning rate over $\{0.01,0.001,0.0001,0.00001\}$.
On HMDA, we train for $30$ epochs. We search over hidden-layer widths of $(100)$, $(128,\operatorname{BN},128)$, $(128,256,128)$, and $(128,\operatorname{BN},256,\operatorname{BN},128)$. We vary batch size over $\{128,256,512\}$ and learning rate over $\{0.001,0.0001,0.00001\}$. We also include a learning rate schedule under which our learning rate is $0.00005$ for the first five epochs and $0.00001$ for the remaining. In addition, we search over weight decays in $\{0,0.0001,0.00001\}$.
On MEPS, we train for $50$ epochs. We search over hidden-layer widths of $(100)$, $(128,\operatorname{BN},128)$, $(128,256,128)$, and $(128,\operatorname{BN},256,\operatorname{BN},128)$. We vary batch size over $\{16,32,64\}$ and learning rate over $\{0.1,0.01,0.01,0.001,0.0001,0.00001\}$. In addition, we search over weight decays in $\{0,0.0001,0.00001\}$.
To ensure all NN outputs are probabilistic, we apply the softmax function before evaluating predictions or passing into any post-processing algorithm. The only exception to this rule is when we apply temperature scaling, which scales raw logits before passing into softmax.
E.2 Models Used on Vision and Language Tasks
On our vision and language datasets, we use much larger models, in some cases with as many as 85 million trainable parameters. We opt for hyperparameters already present in the literature, or alterations of such hyperparameters which give nontrivial accuracy, and we use the same hyperparameters for each calibration fraction. We examine all calibration fractions $\operatorname{CF}\in\{0,0.2,0.4\}$ and average all runs over three random splits of the training and validation data.
Our experiments with language datasets involved two models: (1) DistilBERT, a pretrained transformer introduced by Sanh et al., (2019), and (2) a ResNet-56 using unfrozen, pretrained GloVe embeddings (Pennington et al., , 2014), the implementation for which comes from Duchene et al., (2023).
On the CivilComments dataset, we train a DistilBERT for $10$ epochs with a batch size of $16$, learning rate of $0.00001$, and weight decay of $0.01$, using the Adam optimizer. We fix a maximum token length of $300$.
On the Amazon Polarity dataset, we train a ResNet-56 with three input channels, which accept a stacked embedding of $512$ dimensions. We use the basic_english tokenizer provided by torchtext, fixing a maximum token length of $70$ and minimum frequency of $5$. We train for $10$ epochs with a batch size of $32$ and learning rate of $0.0001$ using Adam. Implementation-specific details are provided in our code.
Our experiments with vision datasets involve three models: (1) vit-large-patch32-224-in21k, a pretrained vision transformer introduced by Dosovitskiy et al., (2021), (2) ResNet-50 (He et al., , 2016), and (3) DenseNet-121 (Huang et al., , 2017).
On CelebA, we train the ViT for $10$ epochs with a batch size fo $64$, learning rate of $0.0001$, and weight decay of $0.01$, using Adam. We also train a ResNet-50 for $50$ epochs with a batch size of $64$ and learning rate of $0.001$, using SGD with a momentum of $0.9$.
On Camelyon17, we train the ViT for $5$ epochs with a batch size of $32$, learning rate of $0.001$, and weight decay of $0.01$. We optimize with SGD, using a momentum of $0.9$. We also train a DenseNet-121 for $10$ epochs with a batch size of $32$, learning rate of $0.001$, and weight decay of $0.01$, using SGD with a momentum of $0.9$.
Appendix F Results on Tabular Datasets
F.1 Plots for All Multicalibration Algorithms
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 20: All multicalibration algorithms on Decision Trees.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 21: All multicalibration algorithms on Random Forest.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 22: All multicalibration algorithms on Logistic Regression.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 23: All multicalibration algorithms on SVMs.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 24: All multicalibration algorithms on Naive Bayes.
F.2 Tables Comparing Best-Performing Multicalibration Algorithms with ERM
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.01 ± 0.003 0.069 ± 0.011 0.012 ± 0.003 0.058 ± 0.005 0.812 ± 0.001 MLP HKRR 0.023 ± 0.001 0.065 ± 0.004 0.023 ± 0.001 0.063 ± 0.002 0.615 ± 0.0 MLP HJZ 0.01 ± 0.002 0.069 ± 0.008 0.013 ± 0.001 0.055 ± 0.004 0.81 ± 0.002 MLP Platt 0.017 ± 0.009 0.076 ± 0.008 0.018 ± 0.008 0.064 ± 0.008 0.809 ± 0.003 MLP Temp 0.011 ± 0.005 0.068 ± 0.01 0.013 ± 0.004 0.059 ± 0.007 0.811 ± 0.001 MLP Isotonic 0.01 ± 0.001 0.067 ± 0.008 0.011 ± 0.001 0.057 ± 0.002 0.811 ± 0.001 RandomForest ERM 0.01 ± 0.001 0.051 ± 0.01 0.011 ± 0.0 0.046 ± 0.002 0.819 ± 0.001 RandomForest HKRR 0.023 ± 0.001 0.066 ± 0.004 0.023 ± 0.001 0.063 ± 0.002 0.614 ± 0.001 RandomForest HJZ 0.007 ± 0.002 0.052 ± 0.003 0.01 ± 0.001 0.047 ± 0.003 0.818 ± 0.001 RandomForest Platt 0.006 ± 0.001 0.054 ± 0.001 0.01 ± 0.001 0.047 ± 0.002 0.818 ± 0.001 RandomForest Temp 0.027 ± 0.001 0.074 ± 0.008 0.027 ± 0.0 0.061 ± 0.004 0.819 ± 0.001 RandomForest Isotonic 0.008 ± 0.001 0.059 ± 0.011 0.011 ± 0.0 0.048 ± 0.004 0.818 ± 0.001 SVM ERM 0.216 ± 0.001 0.268 ± 0.002 0.109 ± 0.0 0.135 ± 0.001 0.784 ± 0.001 SVM HKRR 0.023 ± 0.001 0.065 ± 0.004 0.023 ± 0.001 0.063 ± 0.002 0.615 ± 0.0 SVM HJZ 0.03 ± 0.002 0.074 ± 0.002 0.026 ± 0.002 0.068 ± 0.006 0.612 ± 0.0 SVM Platt 0.336 ± 0.007 0.403 ± 0.003 0.168 ± 0.004 0.2 ± 0.001 0.664 ± 0.007 SVM Temp 0.022 ± 0.005 0.117 ± 0.005 0.022 ± 0.005 0.117 ± 0.005 0.678 ± 0.006 SVM Isotonic 0.081 ± 0.012 0.155 ± 0.007 0.081 ± 0.012 0.146 ± 0.006 0.664 ± 0.007 LogisticRegression ERM 0.012 ± 0.002 0.065 ± 0.011 0.015 ± 0.002 0.063 ± 0.011 0.779 ± 0.007 LogisticRegression HKRR 0.01 ± 0.001 0.042 ± 0.006 0.01 ± 0.001 0.037 ± 0.002 0.783 ± 0.0 LogisticRegression HJZ 0.011 ± 0.001 0.065 ± 0.005 0.014 ± 0.001 0.057 ± 0.002 0.783 ± 0.001 LogisticRegression Platt 0.023 ± 0.006 0.08 ± 0.019 0.024 ± 0.006 0.076 ± 0.02 0.772 ± 0.011 LogisticRegression Temp 0.02 ± 0.001 0.078 ± 0.005 0.021 ± 0.0 0.072 ± 0.003 0.783 ± 0.0 LogisticRegression Isotonic 0.005 ± 0.001 0.068 ± 0.008 0.009 ± 0.001 0.066 ± 0.009 0.775 ± 0.009 DecisionTree ERM 0.017 ± 0.001 0.066 ± 0.01 0.016 ± 0.001 0.059 ± 0.004 0.804 ± 0.0 DecisionTree HKRR 0.023 ± 0.001 0.065 ± 0.004 0.023 ± 0.001 0.063 ± 0.002 0.615 ± 0.0 DecisionTree HJZ 0.013 ± 0.002 0.064 ± 0.005 0.013 ± 0.001 0.054 ± 0.005 0.803 ± 0.002 DecisionTree Platt 0.015 ± 0.002 0.058 ± 0.004 0.014 ± 0.002 0.055 ± 0.004 0.803 ± 0.002 DecisionTree Temp 0.029 ± 0.002 0.088 ± 0.009 0.028 ± 0.002 0.072 ± 0.006 0.803 ± 0.001 DecisionTree Isotonic 0.007 ± 0.002 0.072 ± 0.01 0.01 ± 0.001 0.057 ± 0.003 0.803 ± 0.001 NaiveBayes ERM 0.117 ± 0.0 0.165 ± 0.0 0.109 ± 0.0 0.149 ± 0.001 0.754 ± 0.0 NaiveBayes HKRR 0.023 ± 0.001 0.065 ± 0.004 0.023 ± 0.001 0.063 ± 0.002 0.615 ± 0.0 NaiveBayes HJZ 0.03 ± 0.002 0.074 ± 0.002 0.026 ± 0.002 0.068 ± 0.006 0.612 ± 0.0 NaiveBayes Platt 0.091 ± 0.004 0.13 ± 0.004 0.086 ± 0.004 0.12 ± 0.003 0.759 ± 0.001 NaiveBayes Temp 0.089 ± 0.003 0.154 ± 0.004 0.087 ± 0.002 0.153 ± 0.004 0.754 ± 0.001 NaiveBayes Isotonic 0.004 ± 0.001 0.094 ± 0.003 0.007 ± 0.0 0.085 ± 0.002 0.769 ± 0.001
Figure 25: ACS Income.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.009 ± 0.004 0.048 ± 0.012 0.012 ± 0.002 0.046 ± 0.01 0.901 ± 0.002 MLP HKRR 0.007 ± 0.001 0.044 ± 0.006 0.007 ± 0.002 0.039 ± 0.003 0.879 ± 0.0 MLP HJZ 0.01 ± 0.002 0.043 ± 0.011 0.013 ± 0.002 0.039 ± 0.007 0.9 ± 0.003 MLP Platt 0.01 ± 0.002 0.048 ± 0.012 0.012 ± 0.001 0.044 ± 0.01 0.899 ± 0.001 MLP Temp 0.021 ± 0.006 0.047 ± 0.005 0.022 ± 0.005 0.041 ± 0.003 0.9 ± 0.002 MLP Isotonic 0.014 ± 0.003 0.044 ± 0.009 0.015 ± 0.002 0.04 ± 0.007 0.9 ± 0.0 RandomForest ERM 0.014 ± 0.001 0.045 ± 0.003 0.015 ± 0.0 0.038 ± 0.002 0.903 ± 0.002 RandomForest HKRR 0.007 ± 0.001 0.044 ± 0.006 0.007 ± 0.002 0.039 ± 0.003 0.879 ± 0.0 RandomForest HJZ 0.008 ± 0.001 0.035 ± 0.005 0.011 ± 0.001 0.031 ± 0.003 0.902 ± 0.001 RandomForest Platt 0.01 ± 0.002 0.039 ± 0.002 0.013 ± 0.001 0.033 ± 0.002 0.903 ± 0.001 RandomForest Temp 0.06 ± 0.002 0.084 ± 0.005 0.056 ± 0.001 0.07 ± 0.003 0.899 ± 0.001 RandomForest Isotonic 0.013 ± 0.005 0.037 ± 0.009 0.015 ± 0.004 0.034 ± 0.004 0.902 ± 0.001 SVM ERM 0.106 ± 0.001 0.211 ± 0.019 0.053 ± 0.001 0.106 ± 0.009 0.894 ± 0.001 SVM HKRR 0.007 ± 0.001 0.044 ± 0.006 0.007 ± 0.002 0.039 ± 0.003 0.879 ± 0.0 SVM HJZ 0.003 ± 0.001 0.073 ± 0.005 0.009 ± 0.001 0.073 ± 0.005 0.879 ± 0.0 SVM Platt 0.117 ± 0.001 0.246 ± 0.005 0.059 ± 0.001 0.123 ± 0.003 0.883 ± 0.001 SVM Temp 0.041 ± 0.004 0.091 ± 0.001 0.041 ± 0.004 0.091 ± 0.001 0.879 ± 0.001 SVM Isotonic 0.023 ± 0.009 0.129 ± 0.024 0.023 ± 0.009 0.129 ± 0.023 0.88 ± 0.001 LogisticRegression ERM 0.032 ± 0.001 0.062 ± 0.01 0.03 ± 0.001 0.053 ± 0.002 0.899 ± 0.001 LogisticRegression HKRR 0.007 ± 0.001 0.044 ± 0.006 0.007 ± 0.002 0.039 ± 0.003 0.879 ± 0.0 LogisticRegression HJZ 0.011 ± 0.001 0.045 ± 0.004 0.015 ± 0.001 0.042 ± 0.001 0.9 ± 0.001 LogisticRegression Platt 0.012 ± 0.001 0.049 ± 0.008 0.016 ± 0.001 0.043 ± 0.005 0.899 ± 0.002 LogisticRegression Temp 0.062 ± 0.001 0.088 ± 0.007 0.055 ± 0.001 0.066 ± 0.003 0.899 ± 0.001 LogisticRegression Isotonic 0.009 ± 0.002 0.04 ± 0.006 0.013 ± 0.001 0.036 ± 0.006 0.899 ± 0.002 DecisionTree ERM 0.028 ± 0.002 0.096 ± 0.014 0.022 ± 0.001 0.069 ± 0.003 0.897 ± 0.002 DecisionTree HKRR 0.007 ± 0.001 0.044 ± 0.006 0.007 ± 0.002 0.039 ± 0.003 0.879 ± 0.0 DecisionTree HJZ 0.003 ± 0.001 0.073 ± 0.005 0.009 ± 0.001 0.073 ± 0.005 0.879 ± 0.0 DecisionTree Platt 0.028 ± 0.004 0.086 ± 0.01 0.023 ± 0.003 0.067 ± 0.008 0.897 ± 0.003 DecisionTree Temp 0.056 ± 0.003 0.092 ± 0.01 0.05 ± 0.002 0.073 ± 0.002 0.896 ± 0.002 DecisionTree Isotonic 0.01 ± 0.002 0.06 ± 0.004 0.011 ± 0.002 0.055 ± 0.005 0.896 ± 0.002 NaiveBayes ERM 0.122 ± 0.003 0.271 ± 0.002 0.093 ± 0.002 0.197 ± 0.007 0.857 ± 0.003 NaiveBayes HKRR 0.007 ± 0.001 0.044 ± 0.006 0.007 ± 0.002 0.039 ± 0.003 0.879 ± 0.0 NaiveBayes HJZ 0.003 ± 0.001 0.073 ± 0.005 0.009 ± 0.001 0.073 ± 0.005 0.879 ± 0.0 NaiveBayes Platt 0.121 ± 0.003 0.263 ± 0.006 0.094 ± 0.002 0.195 ± 0.008 0.857 ± 0.002 NaiveBayes Temp 0.083 ± 0.003 0.264 ± 0.008 0.08 ± 0.003 0.24 ± 0.01 0.858 ± 0.002 NaiveBayes Isotonic 0.011 ± 0.003 0.055 ± 0.009 0.014 ± 0.003 0.047 ± 0.006 0.885 ± 0.002
Figure 26: Bank Marketing.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.05 ± 0.007 0.104 ± 0.013 0.049 ± 0.006 0.092 ± 0.013 0.824 ± 0.007 MLP HKRR 0.005 ± 0.001 0.08 ± 0.006 0.005 ± 0.001 0.076 ± 0.005 0.754 ± 0.0 MLP HJZ 0.014 ± 0.003 0.076 ± 0.013 0.016 ± 0.002 0.065 ± 0.007 0.829 ± 0.009 MLP Platt 0.132 ± 0.01 0.176 ± 0.013 0.12 ± 0.007 0.149 ± 0.011 0.816 ± 0.008 MLP Temp 0.022 ± 0.007 0.083 ± 0.014 0.022 ± 0.006 0.078 ± 0.013 0.83 ± 0.007 MLP Isotonic 0.009 ± 0.001 0.076 ± 0.011 0.011 ± 0.001 0.071 ± 0.009 0.831 ± 0.007 RandomForest ERM 0.038 ± 0.002 0.099 ± 0.008 0.038 ± 0.002 0.088 ± 0.006 0.868 ± 0.001 RandomForest HKRR 0.013 ± 0.001 0.061 ± 0.019 0.013 ± 0.001 0.047 ± 0.006 0.862 ± 0.002 RandomForest HJZ 0.024 ± 0.003 0.076 ± 0.01 0.024 ± 0.002 0.062 ± 0.008 0.852 ± 0.022 RandomForest Platt 0.017 ± 0.002 0.078 ± 0.009 0.017 ± 0.002 0.069 ± 0.006 0.868 ± 0.001 RandomForest Temp 0.04 ± 0.002 0.061 ± 0.003 0.04 ± 0.001 0.055 ± 0.004 0.867 ± 0.001 RandomForest Isotonic 0.009 ± 0.002 0.058 ± 0.008 0.01 ± 0.002 0.048 ± 0.004 0.869 ± 0.001 SVM ERM 0.144 ± 0.001 0.175 ± 0.004 0.072 ± 0.0 0.088 ± 0.002 0.856 ± 0.001 SVM HKRR 0.005 ± 0.001 0.08 ± 0.006 0.005 ± 0.001 0.076 ± 0.005 0.754 ± 0.0 SVM HJZ 0.051 ± 0.006 0.133 ± 0.006 0.047 ± 0.009 0.133 ± 0.006 0.721 ± 0.02 SVM Platt 0.353 ± 0.003 0.417 ± 0.006 0.175 ± 0.001 0.205 ± 0.002 0.647 ± 0.003 SVM Temp 0.268 ± 0.001 0.288 ± 0.003 0.254 ± 0.001 0.269 ± 0.002 0.631 ± 0.005 SVM Isotonic 0.06 ± 0.049 0.187 ± 0.033 0.044 ± 0.036 0.152 ± 0.028 0.754 ± 0.0 LogisticRegression ERM 0.016 ± 0.001 0.103 ± 0.002 0.016 ± 0.001 0.101 ± 0.002 0.827 ± 0.001 LogisticRegression HKRR 0.012 ± 0.001 0.043 ± 0.01 0.012 ± 0.0 0.04 ± 0.011 0.83 ± 0.002 LogisticRegression HJZ 0.023 ± 0.006 0.084 ± 0.014 0.024 ± 0.006 0.076 ± 0.019 0.833 ± 0.01 LogisticRegression Platt 0.019 ± 0.007 0.079 ± 0.016 0.02 ± 0.006 0.077 ± 0.016 0.831 ± 0.011 LogisticRegression Temp 0.062 ± 0.012 0.1 ± 0.017 0.062 ± 0.01 0.095 ± 0.015 0.832 ± 0.012 LogisticRegression Isotonic 0.004 ± 0.001 0.088 ± 0.018 0.007 ± 0.001 0.087 ± 0.019 0.832 ± 0.012 DecisionTree ERM 0.019 ± 0.001 0.064 ± 0.007 0.018 ± 0.002 0.055 ± 0.006 0.863 ± 0.001 DecisionTree HKRR 0.013 ± 0.002 0.056 ± 0.014 0.013 ± 0.002 0.051 ± 0.011 0.858 ± 0.001 DecisionTree HJZ 0.018 ± 0.002 0.07 ± 0.013 0.017 ± 0.002 0.058 ± 0.007 0.862 ± 0.001 DecisionTree Platt 0.017 ± 0.002 0.073 ± 0.007 0.016 ± 0.001 0.057 ± 0.009 0.863 ± 0.002 DecisionTree Temp 0.064 ± 0.002 0.09 ± 0.007 0.055 ± 0.001 0.07 ± 0.004 0.859 ± 0.001 DecisionTree Isotonic 0.011 ± 0.004 0.066 ± 0.007 0.013 ± 0.003 0.057 ± 0.008 0.86 ± 0.003 NaiveBayes ERM 0.134 ± 0.001 0.199 ± 0.003 0.126 ± 0.0 0.165 ± 0.002 0.808 ± 0.001 NaiveBayes HKRR 0.009 ± 0.002 0.062 ± 0.008 0.009 ± 0.002 0.059 ± 0.009 0.817 ± 0.003 NaiveBayes HJZ 0.052 ± 0.012 0.117 ± 0.013 0.052 ± 0.01 0.108 ± 0.009 0.805 ± 0.003 NaiveBayes Platt 0.124 ± 0.003 0.2 ± 0.009 0.118 ± 0.003 0.168 ± 0.006 0.809 ± 0.002 NaiveBayes Temp 0.174 ± 0.002 0.185 ± 0.001 0.173 ± 0.002 0.177 ± 0.002 0.809 ± 0.0 NaiveBayes Isotonic 0.006 ± 0.001 0.116 ± 0.002 0.009 ± 0.001 0.116 ± 0.002 0.817 ± 0.001
Figure 27: HMDA.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.018 ± 0.006 0.116 ± 0.035 0.02 ± 0.005 0.061 ± 0.005 0.819 ± 0.001 MLP HKRR 0.029 ± 0.002 0.057 ± 0.005 0.026 ± 0.001 0.049 ± 0.005 0.781 ± 0.0 MLP HJZ 0.029 ± 0.003 0.086 ± 0.003 0.028 ± 0.002 0.073 ± 0.001 0.781 ± 0.0 MLP Platt 0.028 ± 0.002 0.086 ± 0.001 0.027 ± 0.001 0.075 ± 0.002 0.781 ± 0.0 MLP Temp 0.035 ± 0.008 0.098 ± 0.015 0.035 ± 0.008 0.064 ± 0.004 0.821 ± 0.001 MLP Isotonic 0.003 ± 0.001 0.059 ± 0.001 0.003 ± 0.001 0.059 ± 0.001 0.781 ± 0.0 RandomForest ERM 0.019 ± 0.001 0.112 ± 0.017 0.02 ± 0.001 0.051 ± 0.003 0.82 ± 0.001 RandomForest HKRR 0.029 ± 0.002 0.057 ± 0.005 0.026 ± 0.001 0.049 ± 0.005 0.781 ± 0.0 RandomForest HJZ 0.029 ± 0.003 0.086 ± 0.003 0.028 ± 0.002 0.073 ± 0.001 0.781 ± 0.0 RandomForest Platt 0.027 ± 0.006 0.125 ± 0.028 0.029 ± 0.005 0.07 ± 0.008 0.812 ± 0.01 RandomForest Temp 0.027 ± 0.002 0.071 ± 0.006 0.026 ± 0.001 0.048 ± 0.001 0.82 ± 0.002 RandomForest Isotonic 0.027 ± 0.008 0.114 ± 0.015 0.025 ± 0.006 0.059 ± 0.004 0.818 ± 0.001 SVM ERM 0.181 ± 0.0 0.236 ± 0.002 0.091 ± 0.0 0.118 ± 0.001 0.819 ± 0.0 SVM HKRR 0.029 ± 0.002 0.057 ± 0.005 0.026 ± 0.001 0.049 ± 0.005 0.781 ± 0.0 SVM HJZ 0.029 ± 0.003 0.086 ± 0.003 0.028 ± 0.002 0.073 ± 0.001 0.781 ± 0.0 SVM Platt 0.18 ± 0.0 0.232 ± 0.003 0.09 ± 0.0 0.116 ± 0.001 0.82 ± 0.0 SVM Temp 0.06 ± 0.001 0.088 ± 0.0 0.06 ± 0.001 0.088 ± 0.0 0.819 ± 0.0 SVM Isotonic 0.013 ± 0.005 0.04 ± 0.005 0.013 ± 0.005 0.04 ± 0.004 0.82 ± 0.0 LogisticRegression ERM 0.01 ± 0.001 0.102 ± 0.026 0.015 ± 0.001 0.056 ± 0.002 0.819 ± 0.0 LogisticRegression HKRR 0.029 ± 0.002 0.057 ± 0.005 0.026 ± 0.001 0.049 ± 0.005 0.781 ± 0.0 LogisticRegression HJZ 0.029 ± 0.003 0.086 ± 0.003 0.028 ± 0.002 0.073 ± 0.001 0.781 ± 0.0 LogisticRegression Platt 0.023 ± 0.005 0.114 ± 0.018 0.023 ± 0.004 0.068 ± 0.005 0.817 ± 0.002 LogisticRegression Temp 0.022 ± 0.003 0.101 ± 0.015 0.022 ± 0.002 0.056 ± 0.003 0.819 ± 0.001 LogisticRegression Isotonic 0.009 ± 0.002 0.115 ± 0.026 0.014 ± 0.002 0.06 ± 0.004 0.818 ± 0.001 DecisionTree ERM 0.04 ± 0.003 0.181 ± 0.01 0.031 ± 0.001 0.089 ± 0.006 0.81 ± 0.003 DecisionTree HKRR 0.029 ± 0.002 0.057 ± 0.005 0.026 ± 0.001 0.049 ± 0.005 0.781 ± 0.0 DecisionTree HJZ 0.029 ± 0.003 0.086 ± 0.003 0.028 ± 0.002 0.073 ± 0.001 0.781 ± 0.0 DecisionTree Platt 0.038 ± 0.003 0.138 ± 0.027 0.029 ± 0.003 0.08 ± 0.007 0.811 ± 0.003 DecisionTree Temp 0.077 ± 0.001 0.154 ± 0.015 0.075 ± 0.001 0.103 ± 0.004 0.81 ± 0.003 DecisionTree Isotonic 0.021 ± 0.006 0.084 ± 0.018 0.022 ± 0.005 0.07 ± 0.007 0.811 ± 0.005 NaiveBayes ERM 0.187 ± 0.006 0.248 ± 0.009 0.108 ± 0.005 0.137 ± 0.004 0.807 ± 0.004 NaiveBayes HKRR 0.029 ± 0.002 0.057 ± 0.005 0.026 ± 0.001 0.049 ± 0.005 0.781 ± 0.0 NaiveBayes HJZ 0.029 ± 0.003 0.086 ± 0.003 0.028 ± 0.002 0.073 ± 0.001 0.781 ± 0.0 NaiveBayes Platt 0.197 ± 0.011 0.257 ± 0.012 0.119 ± 0.016 0.154 ± 0.019 0.792 ± 0.014 NaiveBayes Temp 0.07 ± 0.019 0.1 ± 0.023 0.069 ± 0.019 0.097 ± 0.025 0.807 ± 0.003 NaiveBayes Isotonic 0.028 ± 0.006 0.09 ± 0.022 0.027 ± 0.005 0.057 ± 0.008 0.806 ± 0.005
Figure 28: Credit Default.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.022 ± 0.006 0.106 ± 0.009 0.024 ± 0.002 0.086 ± 0.015 0.864 ± 0.001 MLP HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 MLP HJZ 0.019 ± 0.003 0.088 ± 0.011 0.021 ± 0.002 0.076 ± 0.018 0.864 ± 0.003 MLP Platt 0.017 ± 0.005 0.1 ± 0.019 0.019 ± 0.003 0.088 ± 0.02 0.865 ± 0.003 MLP Temp 0.019 ± 0.007 0.091 ± 0.016 0.02 ± 0.004 0.081 ± 0.02 0.866 ± 0.001 MLP Isotonic 0.02 ± 0.006 0.108 ± 0.021 0.02 ± 0.004 0.089 ± 0.021 0.864 ± 0.003 RandomForest ERM 0.019 ± 0.001 0.094 ± 0.006 0.021 ± 0.001 0.083 ± 0.004 0.863 ± 0.003 RandomForest HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 RandomForest HJZ 0.021 ± 0.004 0.106 ± 0.011 0.021 ± 0.003 0.101 ± 0.012 0.86 ± 0.003 RandomForest Platt 0.017 ± 0.003 0.093 ± 0.003 0.02 ± 0.001 0.085 ± 0.005 0.861 ± 0.006 RandomForest Temp 0.045 ± 0.003 0.096 ± 0.007 0.045 ± 0.002 0.092 ± 0.009 0.863 ± 0.002 RandomForest Isotonic 0.015 ± 0.002 0.089 ± 0.014 0.017 ± 0.001 0.084 ± 0.014 0.862 ± 0.002 SVM ERM 0.143 ± 0.002 0.376 ± 0.012 0.072 ± 0.001 0.186 ± 0.006 0.857 ± 0.002 SVM HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 SVM HJZ 0.031 ± 0.003 0.156 ± 0.021 0.027 ± 0.004 0.155 ± 0.02 0.828 ± 0.002 SVM Platt 0.14 ± 0.001 0.322 ± 0.019 0.07 ± 0.001 0.161 ± 0.009 0.86 ± 0.001 SVM Temp 0.073 ± 0.008 0.163 ± 0.019 0.073 ± 0.009 0.158 ± 0.015 0.86 ± 0.001 SVM Isotonic 0.048 ± 0.023 0.231 ± 0.085 0.048 ± 0.023 0.218 ± 0.069 0.847 ± 0.017 LogisticRegression ERM 0.022 ± 0.002 0.106 ± 0.008 0.022 ± 0.001 0.083 ± 0.003 0.866 ± 0.002 LogisticRegression HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 LogisticRegression HJZ 0.021 ± 0.003 0.114 ± 0.019 0.023 ± 0.001 0.09 ± 0.011 0.866 ± 0.003 LogisticRegression Platt 0.018 ± 0.003 0.109 ± 0.009 0.021 ± 0.002 0.093 ± 0.017 0.864 ± 0.003 LogisticRegression Temp 0.047 ± 0.002 0.119 ± 0.007 0.044 ± 0.001 0.087 ± 0.003 0.866 ± 0.002 LogisticRegression Isotonic 0.017 ± 0.003 0.109 ± 0.019 0.019 ± 0.003 0.097 ± 0.02 0.863 ± 0.002 DecisionTree ERM 0.067 ± 0.004 0.261 ± 0.028 0.047 ± 0.004 0.166 ± 0.012 0.85 ± 0.006 DecisionTree HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 DecisionTree HJZ 0.031 ± 0.003 0.156 ± 0.021 0.027 ± 0.004 0.155 ± 0.02 0.828 ± 0.002 DecisionTree Platt 0.08 ± 0.005 0.316 ± 0.029 0.054 ± 0.004 0.192 ± 0.009 0.838 ± 0.004 DecisionTree Temp 0.098 ± 0.007 0.214 ± 0.025 0.092 ± 0.005 0.172 ± 0.015 0.838 ± 0.004 DecisionTree Isotonic 0.014 ± 0.003 0.196 ± 0.026 0.015 ± 0.003 0.186 ± 0.027 0.838 ± 0.01 NaiveBayes ERM 0.277 ± 0.019 0.544 ± 0.02 0.164 ± 0.013 0.287 ± 0.011 0.714 ± 0.018 NaiveBayes HKRR 0.019 ± 0.005 0.122 ± 0.008 0.019 ± 0.004 0.104 ± 0.002 0.835 ± 0.003 NaiveBayes HJZ 0.031 ± 0.003 0.156 ± 0.021 0.027 ± 0.004 0.155 ± 0.02 0.828 ± 0.002 NaiveBayes Platt 0.269 ± 0.008 0.535 ± 0.009 0.165 ± 0.004 0.292 ± 0.003 0.719 ± 0.007 NaiveBayes Temp 0.294 ± 0.003 0.368 ± 0.008 0.274 ± 0.002 0.323 ± 0.005 0.719 ± 0.007 NaiveBayes Isotonic 0.019 ± 0.005 0.128 ± 0.017 0.021 ± 0.005 0.122 ± 0.015 0.831 ± 0.006
Figure 29: MEPS.
F.3 Influence of Calibration Fraction on Multicalibration Error and Accuracy
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 30: Influence of calibration fraction on MLP multicalibration and accuracy.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 31: Influence of calibration fraction on decision tree multicalibration.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 32: Influence of calibration fraction on RandomForest multicalibration.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 33: Influence of calibration fraction on LogisticRegression multicalibration.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 34: Influence of calibration fraction on SVM multicalibration.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 35: Influence of calibration fraction on NaiveBayes multicalibration.
F.4 Tables Comparing Multicalibration Algorithms on Reused Data with ERM
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.017 ± 0.003 0.071 ± 0.009 0.017 ± 0.003 0.058 ± 0.009 0.81 ± 0.001 MLP HKRR 0.009 ± 0.002 0.048 ± 0.009 0.009 ± 0.002 0.04 ± 0.006 0.811 ± 0.001 MLP HJZ 0.013 ± 0.003 0.072 ± 0.009 0.015 ± 0.003 0.059 ± 0.007 0.808 ± 0.004 MLP Platt 0.01 ± 0.002 0.074 ± 0.008 0.012 ± 0.001 0.059 ± 0.004 0.811 ± 0.002 MLP Temp 0.017 ± 0.003 0.071 ± 0.009 0.017 ± 0.003 0.058 ± 0.009 0.81 ± 0.001 MLP Isotonic 0.008 ± 0.001 0.07 ± 0.002 0.01 ± 0.001 0.061 ± 0.002 0.811 ± 0.002 RandomForest ERM 0.009 ± 0.001 0.062 ± 0.009 0.011 ± 0.001 0.05 ± 0.003 0.819 ± 0.001 RandomForest HKRR 0.049 ± 0.001 0.123 ± 0.003 0.049 ± 0.001 0.12 ± 0.002 0.819 ± 0.001 RandomForest HJZ 0.007 ± 0.001 0.059 ± 0.01 0.009 ± 0.001 0.048 ± 0.003 0.819 ± 0.001 RandomForest Platt 0.017 ± 0.001 0.058 ± 0.003 0.017 ± 0.001 0.048 ± 0.002 0.819 ± 0.001 RandomForest Temp 0.027 ± 0.001 0.078 ± 0.007 0.027 ± 0.0 0.06 ± 0.001 0.819 ± 0.001 RandomForest Isotonic 0.075 ± 0.0 0.099 ± 0.004 0.071 ± 0.0 0.09 ± 0.003 0.819 ± 0.001 SVM ERM 0.299 ± 0.028 0.38 ± 0.054 0.15 ± 0.013 0.188 ± 0.024 0.701 ± 0.028 SVM HKRR 0.088 ± 0.012 0.061 ± 0.019 0.046 ± 0.006 0.052 ± 0.011 0.704 ± 0.025 SVM HJZ 0.135 ± 0.037 0.197 ± 0.04 0.084 ± 0.032 0.134 ± 0.039 0.703 ± 0.027 SVM Platt 0.299 ± 0.028 0.38 ± 0.054 0.15 ± 0.013 0.188 ± 0.024 0.701 ± 0.028 SVM Temp 0.064 ± 0.034 0.171 ± 0.029 0.062 ± 0.033 0.164 ± 0.031 0.701 ± 0.028 SVM Isotonic 0.002 ± 0.001 0.118 ± 0.014 0.002 ± 0.001 0.118 ± 0.014 0.701 ± 0.028 LogisticRegression ERM 0.012 ± 0.002 0.065 ± 0.011 0.015 ± 0.002 0.063 ± 0.011 0.779 ± 0.007 LogisticRegression HKRR 0.011 ± 0.005 0.045 ± 0.011 0.011 ± 0.005 0.038 ± 0.007 0.781 ± 0.006 LogisticRegression HJZ 0.011 ± 0.004 0.066 ± 0.014 0.013 ± 0.003 0.062 ± 0.015 0.78 ± 0.007 LogisticRegression Platt 0.012 ± 0.004 0.069 ± 0.018 0.013 ± 0.003 0.064 ± 0.016 0.78 ± 0.006 LogisticRegression Temp 0.02 ± 0.001 0.08 ± 0.01 0.02 ± 0.0 0.076 ± 0.009 0.779 ± 0.007 LogisticRegression Isotonic 0.005 ± 0.001 0.064 ± 0.007 0.008 ± 0.001 0.062 ± 0.008 0.779 ± 0.007 DecisionTree ERM 0.017 ± 0.001 0.066 ± 0.01 0.016 ± 0.001 0.059 ± 0.004 0.804 ± 0.0 DecisionTree HKRR 0.016 ± 0.001 0.05 ± 0.004 0.016 ± 0.001 0.047 ± 0.004 0.804 ± 0.0 DecisionTree HJZ 0.017 ± 0.001 0.06 ± 0.007 0.016 ± 0.001 0.055 ± 0.002 0.803 ± 0.001 DecisionTree Platt 0.017 ± 0.001 0.06 ± 0.007 0.016 ± 0.001 0.055 ± 0.002 0.803 ± 0.001 DecisionTree Temp 0.027 ± 0.001 0.078 ± 0.008 0.027 ± 0.001 0.07 ± 0.006 0.804 ± 0.0 DecisionTree Isotonic 0.017 ± 0.001 0.066 ± 0.01 0.016 ± 0.001 0.059 ± 0.004 0.804 ± 0.0 NaiveBayes ERM 0.117 ± 0.0 0.165 ± 0.0 0.109 ± 0.0 0.149 ± 0.001 0.754 ± 0.0 NaiveBayes HKRR 0.039 ± 0.002 0.059 ± 0.014 0.038 ± 0.002 0.047 ± 0.005 0.764 ± 0.002 NaiveBayes HJZ 0.07 ± 0.013 0.111 ± 0.011 0.064 ± 0.01 0.103 ± 0.009 0.751 ± 0.004 NaiveBayes Platt 0.09 ± 0.001 0.127 ± 0.001 0.085 ± 0.0 0.12 ± 0.001 0.76 ± 0.0 NaiveBayes Temp 0.089 ± 0.001 0.155 ± 0.003 0.087 ± 0.0 0.154 ± 0.001 0.754 ± 0.0 NaiveBayes Isotonic 0.003 ± 0.001 0.094 ± 0.002 0.007 ± 0.0 0.086 ± 0.0 0.768 ± 0.0
Figure 36: ACS Income. Training data reused for post-processing.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.013 ± 0.005 0.046 ± 0.005 0.014 ± 0.004 0.042 ± 0.005 0.901 ± 0.001 MLP HKRR 0.009 ± 0.002 0.046 ± 0.002 0.009 ± 0.002 0.041 ± 0.002 0.9 ± 0.001 MLP HJZ 0.007 ± 0.002 0.047 ± 0.006 0.011 ± 0.0 0.039 ± 0.005 0.9 ± 0.002 MLP Platt 0.008 ± 0.002 0.046 ± 0.006 0.011 ± 0.001 0.04 ± 0.005 0.901 ± 0.001 MLP Temp 0.013 ± 0.005 0.046 ± 0.005 0.014 ± 0.004 0.042 ± 0.005 0.901 ± 0.001 MLP Isotonic 0.008 ± 0.002 0.047 ± 0.007 0.01 ± 0.001 0.042 ± 0.006 0.9 ± 0.001 RandomForest ERM 0.014 ± 0.002 0.04 ± 0.003 0.015 ± 0.001 0.037 ± 0.002 0.902 ± 0.001 RandomForest HKRR 0.047 ± 0.003 0.135 ± 0.006 0.046 ± 0.002 0.132 ± 0.006 0.902 ± 0.001 RandomForest HJZ 0.01 ± 0.001 0.038 ± 0.004 0.012 ± 0.001 0.035 ± 0.002 0.903 ± 0.001 RandomForest Platt 0.015 ± 0.001 0.054 ± 0.006 0.017 ± 0.001 0.045 ± 0.002 0.903 ± 0.001 RandomForest Temp 0.058 ± 0.001 0.086 ± 0.003 0.056 ± 0.001 0.076 ± 0.001 0.902 ± 0.001 RandomForest Isotonic 0.056 ± 0.001 0.117 ± 0.005 0.045 ± 0.001 0.093 ± 0.006 0.902 ± 0.001 SVM ERM 0.205 ± 0.11 0.309 ± 0.087 0.102 ± 0.055 0.154 ± 0.041 0.795 ± 0.11 SVM HKRR 0.007 ± 0.001 0.042 ± 0.005 0.007 ± 0.001 0.037 ± 0.002 0.88 ± 0.001 SVM HJZ 0.021 ± 0.005 0.121 ± 0.012 0.024 ± 0.002 0.119 ± 0.014 0.878 ± 0.002 SVM Platt 0.205 ± 0.11 0.309 ± 0.087 0.102 ± 0.055 0.154 ± 0.041 0.795 ± 0.11 SVM Temp 0.165 ± 0.113 0.218 ± 0.105 0.155 ± 0.094 0.205 ± 0.081 0.795 ± 0.11 SVM Isotonic 0.004 ± 0.001 0.144 ± 0.006 0.004 ± 0.001 0.143 ± 0.006 0.879 ± 0.0 LogisticRegression ERM 0.032 ± 0.001 0.062 ± 0.01 0.03 ± 0.001 0.053 ± 0.002 0.899 ± 0.001 LogisticRegression HKRR 0.014 ± 0.003 0.039 ± 0.007 0.014 ± 0.003 0.036 ± 0.006 0.897 ± 0.002 LogisticRegression HJZ 0.01 ± 0.001 0.048 ± 0.005 0.014 ± 0.001 0.04 ± 0.003 0.899 ± 0.001 LogisticRegression Platt 0.012 ± 0.001 0.048 ± 0.006 0.015 ± 0.001 0.04 ± 0.003 0.899 ± 0.001 LogisticRegression Temp 0.062 ± 0.001 0.084 ± 0.004 0.056 ± 0.0 0.064 ± 0.001 0.899 ± 0.001 LogisticRegression Isotonic 0.006 ± 0.002 0.04 ± 0.005 0.01 ± 0.002 0.034 ± 0.003 0.9 ± 0.001 DecisionTree ERM 0.029 ± 0.002 0.099 ± 0.017 0.022 ± 0.001 0.069 ± 0.006 0.897 ± 0.002 DecisionTree HKRR 0.029 ± 0.003 0.101 ± 0.018 0.029 ± 0.003 0.09 ± 0.013 0.897 ± 0.002 DecisionTree HJZ 0.029 ± 0.003 0.1 ± 0.018 0.022 ± 0.002 0.07 ± 0.01 0.897 ± 0.002 DecisionTree Platt 0.028 ± 0.002 0.103 ± 0.017 0.022 ± 0.001 0.071 ± 0.007 0.897 ± 0.002 DecisionTree Temp 0.053 ± 0.004 0.085 ± 0.008 0.048 ± 0.001 0.072 ± 0.003 0.897 ± 0.002 DecisionTree Isotonic 0.029 ± 0.002 0.099 ± 0.017 0.022 ± 0.001 0.069 ± 0.006 0.897 ± 0.002 NaiveBayes ERM 0.122 ± 0.003 0.271 ± 0.002 0.093 ± 0.002 0.197 ± 0.007 0.857 ± 0.003 NaiveBayes HKRR 0.011 ± 0.005 0.044 ± 0.01 0.011 ± 0.005 0.038 ± 0.007 0.88 ± 0.002 NaiveBayes HJZ 0.044 ± 0.009 0.143 ± 0.023 0.042 ± 0.009 0.105 ± 0.011 0.864 ± 0.002 NaiveBayes Platt 0.12 ± 0.003 0.263 ± 0.002 0.094 ± 0.002 0.194 ± 0.007 0.857 ± 0.003 NaiveBayes Temp 0.083 ± 0.002 0.267 ± 0.004 0.081 ± 0.002 0.241 ± 0.005 0.857 ± 0.003 NaiveBayes Isotonic 0.006 ± 0.001 0.047 ± 0.003 0.009 ± 0.001 0.043 ± 0.001 0.886 ± 0.001
Figure 37: Bank Marketing. Training data reused for post-processing.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.034 ± 0.003 0.093 ± 0.014 0.034 ± 0.003 0.08 ± 0.005 0.83 ± 0.005 MLP HKRR 0.011 ± 0.002 0.056 ± 0.015 0.011 ± 0.002 0.05 ± 0.011 0.835 ± 0.004 MLP HJZ 0.009 ± 0.001 0.068 ± 0.01 0.012 ± 0.0 0.058 ± 0.007 0.833 ± 0.004 MLP Platt 0.011 ± 0.002 0.076 ± 0.007 0.013 ± 0.002 0.063 ± 0.003 0.835 ± 0.005 MLP Temp 0.024 ± 0.008 0.086 ± 0.01 0.024 ± 0.008 0.075 ± 0.011 0.83 ± 0.005 MLP Isotonic 0.005 ± 0.001 0.07 ± 0.017 0.008 ± 0.002 0.057 ± 0.005 0.833 ± 0.006 RandomForest ERM 0.038 ± 0.001 0.097 ± 0.007 0.038 ± 0.001 0.089 ± 0.005 0.868 ± 0.001 RandomForest HKRR 0.073 ± 0.002 0.11 ± 0.002 0.073 ± 0.002 0.106 ± 0.002 0.865 ± 0.001 RandomForest HJZ 0.023 ± 0.001 0.058 ± 0.004 0.023 ± 0.001 0.05 ± 0.003 0.868 ± 0.001 RandomForest Platt 0.024 ± 0.001 0.064 ± 0.005 0.024 ± 0.001 0.059 ± 0.004 0.868 ± 0.001 RandomForest Temp 0.038 ± 0.001 0.057 ± 0.003 0.039 ± 0.001 0.051 ± 0.002 0.868 ± 0.001 RandomForest Isotonic 0.074 ± 0.002 0.09 ± 0.005 0.068 ± 0.002 0.079 ± 0.003 0.868 ± 0.001 SVM ERM 0.345 ± 0.2 0.49 ± 0.158 0.166 ± 0.08 0.233 ± 0.055 0.655 ± 0.2 SVM HKRR 0.008 ± 0.002 0.071 ± 0.012 0.008 ± 0.002 0.058 ± 0.012 0.754 ± 0.0 SVM HJZ 0.051 ± 0.017 0.177 ± 0.033 0.048 ± 0.016 0.169 ± 0.03 0.723 ± 0.029 SVM Platt 0.345 ± 0.2 0.49 ± 0.158 0.166 ± 0.08 0.233 ± 0.055 0.655 ± 0.2 SVM Temp 0.135 ± 0.155 0.197 ± 0.155 0.119 ± 0.124 0.176 ± 0.112 0.655 ± 0.2 SVM Isotonic 0.002 ± 0.001 0.163 ± 0.002 0.002 ± 0.001 0.164 ± 0.002 0.754 ± 0.0 LogisticRegression ERM 0.016 ± 0.001 0.103 ± 0.002 0.016 ± 0.001 0.101 ± 0.002 0.827 ± 0.001 LogisticRegression HKRR 0.008 ± 0.001 0.036 ± 0.004 0.008 ± 0.001 0.035 ± 0.003 0.832 ± 0.001 LogisticRegression HJZ 0.017 ± 0.002 0.085 ± 0.006 0.018 ± 0.002 0.083 ± 0.005 0.828 ± 0.001 LogisticRegression Platt 0.008 ± 0.001 0.082 ± 0.004 0.01 ± 0.0 0.082 ± 0.004 0.829 ± 0.001 LogisticRegression Temp 0.069 ± 0.001 0.109 ± 0.002 0.067 ± 0.001 0.102 ± 0.003 0.827 ± 0.001 LogisticRegression Isotonic 0.003 ± 0.0 0.097 ± 0.002 0.006 ± 0.001 0.095 ± 0.001 0.826 ± 0.001 DecisionTree ERM 0.019 ± 0.002 0.064 ± 0.008 0.018 ± 0.002 0.054 ± 0.004 0.863 ± 0.001 DecisionTree HKRR 0.019 ± 0.002 0.052 ± 0.007 0.019 ± 0.002 0.048 ± 0.007 0.864 ± 0.001 DecisionTree HJZ 0.02 ± 0.001 0.058 ± 0.005 0.019 ± 0.001 0.052 ± 0.004 0.863 ± 0.001 DecisionTree Platt 0.019 ± 0.002 0.066 ± 0.006 0.017 ± 0.001 0.055 ± 0.006 0.863 ± 0.001 DecisionTree Temp 0.062 ± 0.002 0.089 ± 0.005 0.054 ± 0.002 0.073 ± 0.003 0.863 ± 0.001 DecisionTree Isotonic 0.019 ± 0.002 0.064 ± 0.008 0.018 ± 0.002 0.054 ± 0.004 0.863 ± 0.001 NaiveBayes ERM 0.134 ± 0.001 0.199 ± 0.003 0.126 ± 0.0 0.165 ± 0.002 0.808 ± 0.001 NaiveBayes HKRR 0.009 ± 0.001 0.069 ± 0.008 0.009 ± 0.001 0.059 ± 0.005 0.814 ± 0.002 NaiveBayes HJZ 0.054 ± 0.009 0.12 ± 0.018 0.052 ± 0.008 0.115 ± 0.019 0.807 ± 0.001 NaiveBayes Platt 0.122 ± 0.0 0.193 ± 0.003 0.116 ± 0.0 0.165 ± 0.002 0.808 ± 0.001 NaiveBayes Temp 0.175 ± 0.001 0.185 ± 0.001 0.174 ± 0.001 0.178 ± 0.001 0.808 ± 0.001 NaiveBayes Isotonic 0.006 ± 0.0 0.117 ± 0.001 0.008 ± 0.0 0.117 ± 0.001 0.817 ± 0.001
Figure 38: HMDA. Training data reused for post-processing.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.018 ± 0.005 0.115 ± 0.043 0.02 ± 0.004 0.064 ± 0.007 0.819 ± 0.001 MLP HKRR 0.016 ± 0.003 0.121 ± 0.052 0.016 ± 0.003 0.083 ± 0.02 0.819 ± 0.001 MLP HJZ 0.017 ± 0.003 0.097 ± 0.011 0.019 ± 0.002 0.06 ± 0.006 0.819 ± 0.002 MLP Platt 0.016 ± 0.002 0.116 ± 0.036 0.018 ± 0.001 0.058 ± 0.007 0.819 ± 0.002 MLP Temp 0.018 ± 0.005 0.115 ± 0.043 0.02 ± 0.004 0.064 ± 0.007 0.819 ± 0.001 MLP Isotonic 0.019 ± 0.002 0.156 ± 0.026 0.018 ± 0.001 0.071 ± 0.005 0.819 ± 0.001 RandomForest ERM 0.019 ± 0.002 0.112 ± 0.021 0.02 ± 0.001 0.052 ± 0.004 0.82 ± 0.001 RandomForest HKRR 0.022 ± 0.003 0.144 ± 0.023 0.022 ± 0.002 0.089 ± 0.008 0.818 ± 0.001 RandomForest HJZ 0.018 ± 0.003 0.123 ± 0.027 0.019 ± 0.002 0.058 ± 0.003 0.818 ± 0.002 RandomForest Platt 0.027 ± 0.003 0.141 ± 0.014 0.026 ± 0.002 0.063 ± 0.002 0.82 ± 0.001 RandomForest Temp 0.026 ± 0.003 0.073 ± 0.01 0.026 ± 0.001 0.048 ± 0.002 0.82 ± 0.001 RandomForest Isotonic 0.037 ± 0.003 0.124 ± 0.02 0.035 ± 0.002 0.073 ± 0.003 0.82 ± 0.001 SVM ERM 0.18 ± 0.0 0.233 ± 0.0 0.09 ± 0.0 0.117 ± 0.0 0.82 ± 0.0 SVM HKRR 0.03 ± 0.001 0.06 ± 0.002 0.021 ± 0.002 0.052 ± 0.004 0.82 ± 0.0 SVM HJZ 0.051 ± 0.006 0.166 ± 0.015 0.038 ± 0.004 0.091 ± 0.006 0.813 ± 0.008 SVM Platt 0.18 ± 0.0 0.233 ± 0.0 0.09 ± 0.0 0.117 ± 0.0 0.82 ± 0.0 SVM Temp 0.057 ± 0.0 0.088 ± 0.0 0.057 ± 0.0 0.088 ± 0.0 0.82 ± 0.0 SVM Isotonic 0.002 ± 0.001 0.044 ± 0.001 0.002 ± 0.001 0.043 ± 0.001 0.82 ± 0.0 LogisticRegression ERM 0.01 ± 0.001 0.102 ± 0.026 0.015 ± 0.001 0.056 ± 0.002 0.819 ± 0.0 LogisticRegression HKRR 0.008 ± 0.001 0.091 ± 0.036 0.008 ± 0.001 0.073 ± 0.028 0.819 ± 0.001 LogisticRegression HJZ 0.013 ± 0.004 0.102 ± 0.032 0.017 ± 0.002 0.059 ± 0.005 0.817 ± 0.002 LogisticRegression Platt 0.013 ± 0.001 0.097 ± 0.035 0.016 ± 0.001 0.057 ± 0.004 0.819 ± 0.0 LogisticRegression Temp 0.023 ± 0.001 0.075 ± 0.011 0.024 ± 0.0 0.053 ± 0.002 0.819 ± 0.0 LogisticRegression Isotonic 0.013 ± 0.003 0.13 ± 0.021 0.017 ± 0.002 0.061 ± 0.005 0.819 ± 0.001 DecisionTree ERM 0.041 ± 0.004 0.186 ± 0.014 0.031 ± 0.002 0.088 ± 0.007 0.81 ± 0.003 DecisionTree HKRR 0.041 ± 0.004 0.183 ± 0.017 0.039 ± 0.003 0.107 ± 0.014 0.81 ± 0.002 DecisionTree HJZ 0.042 ± 0.003 0.168 ± 0.021 0.031 ± 0.001 0.085 ± 0.004 0.81 ± 0.003 DecisionTree Platt 0.043 ± 0.003 0.188 ± 0.02 0.033 ± 0.001 0.089 ± 0.006 0.81 ± 0.003 DecisionTree Temp 0.08 ± 0.001 0.167 ± 0.022 0.076 ± 0.001 0.111 ± 0.014 0.81 ± 0.003 DecisionTree Isotonic 0.041 ± 0.004 0.186 ± 0.014 0.031 ± 0.002 0.088 ± 0.007 0.81 ± 0.003 NaiveBayes ERM 0.187 ± 0.006 0.248 ± 0.009 0.108 ± 0.005 0.137 ± 0.004 0.807 ± 0.004 NaiveBayes HKRR 0.028 ± 0.004 0.091 ± 0.016 0.028 ± 0.004 0.089 ± 0.019 0.807 ± 0.004 NaiveBayes HJZ 0.069 ± 0.018 0.139 ± 0.027 0.057 ± 0.016 0.103 ± 0.018 0.807 ± 0.004 NaiveBayes Platt 0.184 ± 0.007 0.245 ± 0.011 0.11 ± 0.005 0.142 ± 0.005 0.807 ± 0.004 NaiveBayes Temp 0.07 ± 0.02 0.101 ± 0.024 0.068 ± 0.02 0.098 ± 0.026 0.807 ± 0.004 NaiveBayes Isotonic 0.008 ± 0.002 0.076 ± 0.017 0.012 ± 0.001 0.049 ± 0.002 0.81 ± 0.001
Figure 39: Credit Default. Training data reused for post-processing.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.024 ± 0.006 0.107 ± 0.018 0.026 ± 0.004 0.1 ± 0.025 0.865 ± 0.002 MLP HKRR 0.024 ± 0.006 0.109 ± 0.022 0.024 ± 0.006 0.096 ± 0.016 0.862 ± 0.003 MLP HJZ 0.018 ± 0.001 0.105 ± 0.011 0.021 ± 0.002 0.093 ± 0.021 0.864 ± 0.003 MLP Platt 0.019 ± 0.003 0.096 ± 0.016 0.02 ± 0.002 0.084 ± 0.017 0.865 ± 0.002 MLP Temp 0.024 ± 0.006 0.107 ± 0.018 0.026 ± 0.004 0.1 ± 0.025 0.865 ± 0.002 MLP Isotonic 0.017 ± 0.004 0.081 ± 0.008 0.019 ± 0.002 0.07 ± 0.009 0.863 ± 0.002 RandomForest ERM 0.017 ± 0.001 0.091 ± 0.005 0.02 ± 0.001 0.082 ± 0.005 0.862 ± 0.002 RandomForest HKRR 0.089 ± 0.004 0.25 ± 0.026 0.088 ± 0.004 0.221 ± 0.025 0.848 ± 0.002 RandomForest HJZ 0.022 ± 0.003 0.088 ± 0.007 0.024 ± 0.001 0.083 ± 0.004 0.862 ± 0.002 RandomForest Platt 0.027 ± 0.002 0.09 ± 0.003 0.029 ± 0.001 0.08 ± 0.006 0.863 ± 0.001 RandomForest Temp 0.044 ± 0.001 0.102 ± 0.01 0.043 ± 0.001 0.089 ± 0.003 0.862 ± 0.002 RandomForest Isotonic 0.075 ± 0.002 0.167 ± 0.005 0.058 ± 0.002 0.121 ± 0.007 0.86 ± 0.003 SVM ERM 0.149 ± 0.015 0.359 ± 0.04 0.075 ± 0.008 0.179 ± 0.019 0.851 ± 0.015 SVM HKRR 0.018 ± 0.005 0.111 ± 0.013 0.018 ± 0.005 0.101 ± 0.006 0.853 ± 0.01 SVM HJZ 0.06 ± 0.007 0.211 ± 0.069 0.048 ± 0.009 0.183 ± 0.055 0.857 ± 0.006 SVM Platt 0.149 ± 0.015 0.359 ± 0.04 0.075 ± 0.008 0.179 ± 0.019 0.851 ± 0.015 SVM Temp 0.111 ± 0.039 0.198 ± 0.033 0.11 ± 0.038 0.194 ± 0.03 0.851 ± 0.015 SVM Isotonic 0.007 ± 0.004 0.202 ± 0.057 0.007 ± 0.004 0.195 ± 0.051 0.852 ± 0.014 LogisticRegression ERM 0.022 ± 0.002 0.106 ± 0.008 0.022 ± 0.001 0.083 ± 0.003 0.866 ± 0.002 LogisticRegression HKRR 0.024 ± 0.001 0.12 ± 0.011 0.023 ± 0.001 0.105 ± 0.009 0.861 ± 0.003 LogisticRegression HJZ 0.022 ± 0.003 0.106 ± 0.011 0.022 ± 0.002 0.083 ± 0.005 0.866 ± 0.001 LogisticRegression Platt 0.016 ± 0.004 0.092 ± 0.009 0.021 ± 0.002 0.078 ± 0.003 0.863 ± 0.001 LogisticRegression Temp 0.049 ± 0.002 0.112 ± 0.008 0.046 ± 0.001 0.087 ± 0.003 0.866 ± 0.002 LogisticRegression Isotonic 0.014 ± 0.002 0.087 ± 0.01 0.018 ± 0.002 0.073 ± 0.002 0.866 ± 0.001 DecisionTree ERM 0.067 ± 0.006 0.266 ± 0.029 0.048 ± 0.004 0.17 ± 0.013 0.849 ± 0.007 DecisionTree HKRR 0.074 ± 0.007 0.282 ± 0.034 0.073 ± 0.007 0.237 ± 0.022 0.845 ± 0.007 DecisionTree HJZ 0.068 ± 0.007 0.264 ± 0.035 0.049 ± 0.005 0.167 ± 0.018 0.849 ± 0.007 DecisionTree Platt 0.069 ± 0.008 0.267 ± 0.03 0.049 ± 0.005 0.171 ± 0.013 0.849 ± 0.007 DecisionTree Temp 0.095 ± 0.004 0.175 ± 0.019 0.091 ± 0.001 0.155 ± 0.01 0.849 ± 0.007 DecisionTree Isotonic 0.067 ± 0.006 0.266 ± 0.029 0.048 ± 0.004 0.17 ± 0.013 0.849 ± 0.007 NaiveBayes ERM 0.277 ± 0.019 0.544 ± 0.02 0.164 ± 0.013 0.287 ± 0.011 0.714 ± 0.018 NaiveBayes HKRR 0.023 ± 0.003 0.119 ± 0.026 0.023 ± 0.003 0.102 ± 0.019 0.833 ± 0.005 NaiveBayes HJZ 0.05 ± 0.017 0.205 ± 0.038 0.044 ± 0.014 0.183 ± 0.038 0.803 ± 0.004 NaiveBayes Platt 0.275 ± 0.018 0.54 ± 0.019 0.169 ± 0.011 0.295 ± 0.009 0.714 ± 0.018 NaiveBayes Temp 0.3 ± 0.007 0.373 ± 0.016 0.278 ± 0.005 0.326 ± 0.01 0.714 ± 0.018 NaiveBayes Isotonic 0.014 ± 0.002 0.121 ± 0.009 0.017 ± 0.001 0.111 ± 0.005 0.834 ± 0.006
Figure 40: MEPS. Training data reused for post-processing.
F.4.1 Comparing Multicalibration Post-Processing Performance with Data Reuse
Figure 41: Data reuse comparison for ACSIncome.
Figure 42: Data reuse comparison for BankMarketing.
Figure 43: Data reuse comparison for MEPS.
Figure 44: Data reuse comparison for CreditDefault.
Figure 45: Data reuse comparison for HMDA.
Appendix G Results on Tabular Datasets with Alternate Groups
G.1 Plots for All Multicalibration Algorithms
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 46: All multicalibration algorithms on MLPs. Alternate groups.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 47: All multicalibration algorithms on Decision Trees. Alternate groups.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 48: All multicalibration algorithms on Random Forest. Alternate groups.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 49: All multicalibration algorithms on Logistic Regression. Alternate groups.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 50: All multicalibration algorithms on SVMs. Alternate groups.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 51: All multicalibration algorithms on Naive Bayes. Alternate groups.
G.2 Tables Comparing Best-Performing Multicalibration Algorithms with ERM (Alternate Groups)
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.014 ± 0.005 0.059 ± 0.007 0.015 ± 0.004 0.059 ± 0.008 0.811 ± 0.002 MLP HKRR 0.009 ± 0.003 0.037 ± 0.005 0.009 ± 0.003 0.035 ± 0.003 0.803 ± 0.001 MLP HJZ 0.019 ± 0.003 0.053 ± 0.01 0.02 ± 0.002 0.051 ± 0.011 0.807 ± 0.004 MLP Platt 0.012 ± 0.004 0.046 ± 0.008 0.014 ± 0.002 0.044 ± 0.007 0.811 ± 0.001 MLP Temp 0.012 ± 0.005 0.056 ± 0.005 0.013 ± 0.004 0.056 ± 0.005 0.811 ± 0.001 MLP Isotonic 0.012 ± 0.002 0.055 ± 0.005 0.013 ± 0.001 0.054 ± 0.006 0.811 ± 0.001 RandomForest ERM 0.01 ± 0.001 0.036 ± 0.001 0.011 ± 0.0 0.035 ± 0.001 0.819 ± 0.001 RandomForest HKRR 0.007 ± 0.001 0.043 ± 0.011 0.007 ± 0.001 0.042 ± 0.01 0.818 ± 0.0 RandomForest HJZ 0.007 ± 0.001 0.032 ± 0.005 0.011 ± 0.001 0.031 ± 0.003 0.813 ± 0.007 RandomForest Platt 0.009 ± 0.001 0.031 ± 0.003 0.011 ± 0.001 0.029 ± 0.002 0.816 ± 0.004 RandomForest Temp 0.027 ± 0.001 0.074 ± 0.001 0.027 ± 0.001 0.073 ± 0.0 0.819 ± 0.001 RandomForest Isotonic 0.008 ± 0.001 0.033 ± 0.001 0.011 ± 0.0 0.031 ± 0.001 0.818 ± 0.001 SVM ERM 0.216 ± 0.001 0.292 ± 0.004 0.109 ± 0.0 0.149 ± 0.002 0.784 ± 0.001 SVM HKRR 0.008 ± 0.0 0.034 ± 0.005 0.008 ± 0.0 0.032 ± 0.004 0.667 ± 0.0 SVM HJZ 0.013 ± 0.002 0.164 ± 0.009 0.019 ± 0.001 0.163 ± 0.009 0.665 ± 0.0 SVM Platt 0.336 ± 0.007 0.438 ± 0.0 0.168 ± 0.004 0.214 ± 0.0 0.664 ± 0.007 SVM Temp 0.099 ± 0.006 0.216 ± 0.0 0.099 ± 0.006 0.211 ± 0.0 0.678 ± 0.006 SVM Isotonic 0.081 ± 0.012 0.263 ± 0.014 0.081 ± 0.012 0.25 ± 0.011 0.664 ± 0.007 LogisticRegression ERM 0.012 ± 0.002 0.123 ± 0.014 0.015 ± 0.002 0.116 ± 0.012 0.779 ± 0.007 LogisticRegression HKRR 0.006 ± 0.002 0.041 ± 0.003 0.006 ± 0.002 0.038 ± 0.002 0.78 ± 0.007 LogisticRegression HJZ 0.011 ± 0.001 0.085 ± 0.013 0.014 ± 0.001 0.083 ± 0.012 0.775 ± 0.009 LogisticRegression Platt 0.021 ± 0.004 0.108 ± 0.017 0.021 ± 0.004 0.105 ± 0.015 0.774 ± 0.012 LogisticRegression Temp 0.019 ± 0.0 0.115 ± 0.004 0.02 ± 0.0 0.111 ± 0.005 0.776 ± 0.009 LogisticRegression Isotonic 0.005 ± 0.001 0.109 ± 0.007 0.009 ± 0.001 0.105 ± 0.004 0.775 ± 0.009 DecisionTree ERM 0.017 ± 0.001 0.051 ± 0.004 0.016 ± 0.001 0.048 ± 0.004 0.804 ± 0.0 DecisionTree HKRR 0.008 ± 0.001 0.041 ± 0.004 0.008 ± 0.001 0.039 ± 0.003 0.799 ± 0.001 DecisionTree HJZ 0.019 ± 0.001 0.049 ± 0.002 0.017 ± 0.002 0.046 ± 0.002 0.802 ± 0.001 DecisionTree Platt 0.011 ± 0.001 0.041 ± 0.006 0.013 ± 0.001 0.037 ± 0.005 0.803 ± 0.0 DecisionTree Temp 0.028 ± 0.002 0.073 ± 0.001 0.027 ± 0.002 0.072 ± 0.001 0.803 ± 0.001 DecisionTree Isotonic 0.007 ± 0.002 0.054 ± 0.003 0.01 ± 0.001 0.051 ± 0.003 0.803 ± 0.001 NaiveBayes ERM 0.117 ± 0.0 0.201 ± 0.001 0.109 ± 0.0 0.182 ± 0.0 0.754 ± 0.0 NaiveBayes HKRR 0.006 ± 0.001 0.042 ± 0.005 0.006 ± 0.001 0.039 ± 0.004 0.77 ± 0.001 NaiveBayes HJZ 0.017 ± 0.002 0.093 ± 0.006 0.021 ± 0.001 0.091 ± 0.007 0.762 ± 0.004 NaiveBayes Platt 0.085 ± 0.004 0.165 ± 0.007 0.08 ± 0.004 0.161 ± 0.006 0.756 ± 0.002 NaiveBayes Temp 0.079 ± 0.002 0.182 ± 0.002 0.069 ± 0.001 0.18 ± 0.002 0.754 ± 0.001 NaiveBayes Isotonic 0.008 ± 0.002 0.105 ± 0.001 0.011 ± 0.002 0.103 ± 0.001 0.768 ± 0.0
Figure 52: ACS Income. Alternate groups.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.008 ± 0.003 0.14 ± 0.018 0.012 ± 0.002 0.099 ± 0.009 0.9 ± 0.001 MLP HKRR 0.102 ± 0.003 0.127 ± 0.013 0.098 ± 0.001 0.117 ± 0.016 0.879 ± 0.0 MLP HJZ 0.019 ± 0.012 0.137 ± 0.016 0.021 ± 0.011 0.09 ± 0.008 0.901 ± 0.001 MLP Platt 0.011 ± 0.004 0.138 ± 0.022 0.014 ± 0.003 0.105 ± 0.017 0.896 ± 0.005 MLP Temp 0.042 ± 0.01 0.123 ± 0.024 0.042 ± 0.01 0.074 ± 0.007 0.901 ± 0.001 MLP Isotonic 0.008 ± 0.002 0.126 ± 0.024 0.01 ± 0.001 0.091 ± 0.009 0.9 ± 0.001 RandomForest ERM 0.014 ± 0.001 0.095 ± 0.01 0.015 ± 0.0 0.062 ± 0.001 0.903 ± 0.002 RandomForest HKRR 0.012 ± 0.002 0.106 ± 0.017 0.012 ± 0.002 0.082 ± 0.008 0.898 ± 0.001 RandomForest HJZ 0.011 ± 0.003 0.108 ± 0.022 0.013 ± 0.002 0.066 ± 0.014 0.903 ± 0.001 RandomForest Platt 0.009 ± 0.002 0.095 ± 0.022 0.012 ± 0.001 0.06 ± 0.006 0.903 ± 0.001 RandomForest Temp 0.057 ± 0.001 0.117 ± 0.019 0.054 ± 0.001 0.083 ± 0.002 0.903 ± 0.001 RandomForest Isotonic 0.008 ± 0.002 0.116 ± 0.024 0.011 ± 0.001 0.083 ± 0.009 0.901 ± 0.001 SVM ERM 0.106 ± 0.001 0.347 ± 0.027 0.053 ± 0.001 0.173 ± 0.013 0.894 ± 0.001 SVM HKRR 0.051 ± 0.014 0.108 ± 0.038 0.051 ± 0.014 0.091 ± 0.028 0.879 ± 0.0 SVM HJZ 0.106 ± 0.001 0.134 ± 0.003 0.098 ± 0.0 0.112 ± 0.004 0.879 ± 0.0 SVM Platt 0.117 ± 0.001 0.36 ± 0.009 0.059 ± 0.001 0.178 ± 0.004 0.883 ± 0.001 SVM Temp 0.152 ± 0.003 0.225 ± 0.0 0.151 ± 0.003 0.219 ± 0.0 0.88 ± 0.001 SVM Isotonic 0.023 ± 0.009 0.247 ± 0.025 0.023 ± 0.009 0.237 ± 0.021 0.88 ± 0.001 LogisticRegression ERM 0.032 ± 0.001 0.154 ± 0.006 0.03 ± 0.001 0.12 ± 0.002 0.899 ± 0.001 LogisticRegression HKRR 0.018 ± 0.002 0.132 ± 0.027 0.018 ± 0.002 0.108 ± 0.027 0.897 ± 0.002 LogisticRegression HJZ 0.106 ± 0.001 0.134 ± 0.003 0.098 ± 0.0 0.112 ± 0.004 0.879 ± 0.0 LogisticRegression Platt 0.023 ± 0.005 0.156 ± 0.009 0.023 ± 0.004 0.129 ± 0.012 0.897 ± 0.002 LogisticRegression Temp 0.061 ± 0.001 0.174 ± 0.02 0.056 ± 0.0 0.121 ± 0.005 0.899 ± 0.001 LogisticRegression Isotonic 0.008 ± 0.001 0.114 ± 0.012 0.011 ± 0.002 0.093 ± 0.006 0.9 ± 0.001 DecisionTree ERM 0.028 ± 0.002 0.213 ± 0.021 0.022 ± 0.001 0.154 ± 0.016 0.897 ± 0.002 DecisionTree HKRR 0.102 ± 0.003 0.127 ± 0.013 0.098 ± 0.001 0.117 ± 0.016 0.879 ± 0.0 DecisionTree HJZ 0.106 ± 0.001 0.134 ± 0.003 0.098 ± 0.0 0.112 ± 0.004 0.879 ± 0.0 DecisionTree Platt 0.025 ± 0.006 0.214 ± 0.044 0.021 ± 0.002 0.153 ± 0.019 0.897 ± 0.002 DecisionTree Temp 0.116 ± 0.001 0.173 ± 0.013 0.114 ± 0.001 0.163 ± 0.003 0.896 ± 0.002 DecisionTree Isotonic 0.01 ± 0.002 0.157 ± 0.011 0.011 ± 0.002 0.139 ± 0.014 0.896 ± 0.002 NaiveBayes ERM 0.122 ± 0.003 0.521 ± 0.002 0.093 ± 0.002 0.308 ± 0.002 0.857 ± 0.003 NaiveBayes HKRR 0.037 ± 0.005 0.121 ± 0.024 0.036 ± 0.004 0.106 ± 0.016 0.872 ± 0.003 NaiveBayes HJZ 0.106 ± 0.001 0.134 ± 0.003 0.098 ± 0.0 0.112 ± 0.004 0.879 ± 0.0 NaiveBayes Platt 0.122 ± 0.005 0.528 ± 0.01 0.094 ± 0.004 0.318 ± 0.012 0.857 ± 0.004 NaiveBayes Temp 0.217 ± 0.002 0.293 ± 0.009 0.212 ± 0.002 0.268 ± 0.009 0.857 ± 0.004 NaiveBayes Isotonic 0.007 ± 0.001 0.118 ± 0.011 0.01 ± 0.001 0.095 ± 0.012 0.885 ± 0.002
Figure 53: Bank Marketing. Alternate groups.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.018 ± 0.005 0.039 ± 0.004 0.019 ± 0.004 0.035 ± 0.003 0.818 ± 0.001 MLP HKRR 0.028 ± 0.003 0.026 ± 0.003 0.026 ± 0.002 0.025 ± 0.002 0.781 ± 0.0 MLP HJZ 0.033 ± 0.002 0.036 ± 0.008 0.03 ± 0.001 0.028 ± 0.002 0.781 ± 0.0 MLP Platt 0.013 ± 0.004 0.043 ± 0.008 0.015 ± 0.002 0.038 ± 0.008 0.82 ± 0.0 MLP Temp 0.016 ± 0.004 0.043 ± 0.007 0.018 ± 0.003 0.033 ± 0.002 0.819 ± 0.001 MLP Isotonic 0.011 ± 0.003 0.043 ± 0.005 0.014 ± 0.001 0.035 ± 0.004 0.818 ± 0.001 RandomForest ERM 0.019 ± 0.001 0.035 ± 0.003 0.02 ± 0.001 0.033 ± 0.001 0.82 ± 0.001 RandomForest HKRR 0.028 ± 0.003 0.026 ± 0.003 0.026 ± 0.002 0.025 ± 0.002 0.781 ± 0.0 RandomForest HJZ 0.033 ± 0.002 0.036 ± 0.008 0.03 ± 0.001 0.028 ± 0.002 0.781 ± 0.0 RandomForest Platt 0.013 ± 0.002 0.04 ± 0.007 0.016 ± 0.002 0.035 ± 0.003 0.819 ± 0.001 RandomForest Temp 0.023 ± 0.003 0.04 ± 0.003 0.024 ± 0.002 0.037 ± 0.003 0.819 ± 0.001 RandomForest Isotonic 0.013 ± 0.003 0.035 ± 0.002 0.015 ± 0.002 0.031 ± 0.002 0.819 ± 0.001 SVM ERM 0.181 ± 0.0 0.21 ± 0.001 0.091 ± 0.0 0.105 ± 0.0 0.819 ± 0.0 SVM HKRR 0.02 ± 0.001 0.046 ± 0.004 0.016 ± 0.001 0.031 ± 0.003 0.819 ± 0.001 SVM HJZ 0.033 ± 0.002 0.036 ± 0.008 0.03 ± 0.001 0.028 ± 0.002 0.781 ± 0.0 SVM Platt 0.18 ± 0.001 0.213 ± 0.002 0.09 ± 0.0 0.106 ± 0.001 0.82 ± 0.001 SVM Temp 0.022 ± 0.0 0.052 ± 0.001 0.022 ± 0.0 0.051 ± 0.001 0.82 ± 0.0 SVM Isotonic 0.006 ± 0.001 0.026 ± 0.002 0.006 ± 0.001 0.026 ± 0.002 0.82 ± 0.001 LogisticRegression ERM 0.01 ± 0.001 0.042 ± 0.004 0.015 ± 0.001 0.036 ± 0.003 0.819 ± 0.0 LogisticRegression HKRR 0.028 ± 0.003 0.026 ± 0.003 0.026 ± 0.002 0.025 ± 0.002 0.781 ± 0.0 LogisticRegression HJZ 0.033 ± 0.002 0.036 ± 0.008 0.03 ± 0.001 0.028 ± 0.002 0.781 ± 0.0 LogisticRegression Platt 0.014 ± 0.001 0.045 ± 0.004 0.016 ± 0.002 0.037 ± 0.003 0.819 ± 0.0 LogisticRegression Temp 0.023 ± 0.001 0.041 ± 0.006 0.023 ± 0.001 0.037 ± 0.001 0.819 ± 0.0 LogisticRegression Isotonic 0.01 ± 0.002 0.04 ± 0.009 0.014 ± 0.001 0.034 ± 0.006 0.82 ± 0.001 DecisionTree ERM 0.04 ± 0.003 0.078 ± 0.011 0.031 ± 0.001 0.054 ± 0.006 0.81 ± 0.003 DecisionTree HKRR 0.028 ± 0.003 0.026 ± 0.003 0.026 ± 0.002 0.025 ± 0.002 0.781 ± 0.0 DecisionTree HJZ 0.033 ± 0.002 0.036 ± 0.008 0.03 ± 0.001 0.028 ± 0.002 0.781 ± 0.0 DecisionTree Platt 0.033 ± 0.003 0.062 ± 0.005 0.028 ± 0.002 0.05 ± 0.006 0.81 ± 0.003 DecisionTree Temp 0.031 ± 0.004 0.062 ± 0.009 0.029 ± 0.003 0.058 ± 0.01 0.811 ± 0.003 DecisionTree Isotonic 0.007 ± 0.001 0.04 ± 0.006 0.009 ± 0.003 0.036 ± 0.007 0.797 ± 0.004 NaiveBayes ERM 0.187 ± 0.006 0.226 ± 0.005 0.108 ± 0.005 0.132 ± 0.007 0.807 ± 0.004 NaiveBayes HKRR 0.028 ± 0.013 0.035 ± 0.008 0.026 ± 0.013 0.03 ± 0.006 0.784 ± 0.02 NaiveBayes HJZ 0.033 ± 0.002 0.036 ± 0.008 0.03 ± 0.001 0.028 ± 0.002 0.781 ± 0.0 NaiveBayes Platt 0.197 ± 0.011 0.236 ± 0.013 0.119 ± 0.015 0.143 ± 0.023 0.792 ± 0.014 NaiveBayes Temp 0.037 ± 0.012 0.075 ± 0.007 0.037 ± 0.012 0.073 ± 0.008 0.809 ± 0.003 NaiveBayes Isotonic 0.013 ± 0.004 0.039 ± 0.007 0.016 ± 0.003 0.037 ± 0.005 0.809 ± 0.001
Figure 54: Credit Default. Alternate groups.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.045 ± 0.015 0.087 ± 0.018 0.043 ± 0.014 0.084 ± 0.018 0.834 ± 0.002 MLP HKRR 0.004 ± 0.001 0.034 ± 0.006 0.004 ± 0.001 0.033 ± 0.006 0.756 ± 0.0 MLP HJZ 0.012 ± 0.001 0.062 ± 0.021 0.016 ± 0.001 0.056 ± 0.019 0.835 ± 0.009 MLP Platt 0.184 ± 0.009 0.272 ± 0.025 0.14 ± 0.002 0.206 ± 0.017 0.781 ± 0.002 MLP Temp 0.047 ± 0.033 0.097 ± 0.02 0.046 ± 0.032 0.094 ± 0.02 0.817 ± 0.008 MLP Isotonic 0.013 ± 0.002 0.1 ± 0.009 0.014 ± 0.002 0.092 ± 0.008 0.824 ± 0.006 RandomForest ERM 0.038 ± 0.002 0.054 ± 0.001 0.038 ± 0.002 0.053 ± 0.001 0.868 ± 0.001 RandomForest HKRR 0.006 ± 0.001 0.028 ± 0.003 0.006 ± 0.001 0.027 ± 0.003 0.866 ± 0.001 RandomForest HJZ 0.007 ± 0.002 0.044 ± 0.008 0.011 ± 0.002 0.041 ± 0.007 0.869 ± 0.001 RandomForest Platt 0.009 ± 0.001 0.06 ± 0.007 0.011 ± 0.001 0.054 ± 0.005 0.867 ± 0.001 RandomForest Temp 0.037 ± 0.001 0.072 ± 0.004 0.039 ± 0.001 0.07 ± 0.004 0.866 ± 0.001 RandomForest Isotonic 0.009 ± 0.002 0.048 ± 0.002 0.01 ± 0.002 0.046 ± 0.003 0.869 ± 0.001 SVM ERM 0.144 ± 0.001 0.31 ± 0.004 0.072 ± 0.0 0.158 ± 0.002 0.856 ± 0.001 SVM HKRR 0.004 ± 0.001 0.034 ± 0.006 0.004 ± 0.001 0.033 ± 0.006 0.756 ± 0.0 SVM HJZ 0.008 ± 0.002 0.058 ± 0.009 0.011 ± 0.001 0.058 ± 0.009 0.754 ± 0.0 SVM Platt 0.008 ± 0.002 0.079 ± 0.002 0.012 ± 0.001 0.079 ± 0.001 0.754 ± 0.0 SVM Temp 0.266 ± 0.001 0.324 ± 0.005 0.253 ± 0.001 0.295 ± 0.003 0.647 ± 0.003 SVM Isotonic 0.002 ± 0.001 0.115 ± 0.001 0.002 ± 0.001 0.115 ± 0.001 0.754 ± 0.0 LogisticRegression ERM 0.016 ± 0.001 0.171 ± 0.002 0.016 ± 0.001 0.168 ± 0.002 0.827 ± 0.001 LogisticRegression HKRR 0.007 ± 0.002 0.028 ± 0.005 0.007 ± 0.002 0.026 ± 0.005 0.862 ± 0.0 LogisticRegression HJZ 0.006 ± 0.002 0.039 ± 0.007 0.011 ± 0.002 0.037 ± 0.006 0.857 ± 0.003 LogisticRegression Platt 0.008 ± 0.002 0.079 ± 0.002 0.012 ± 0.001 0.079 ± 0.001 0.754 ± 0.0 LogisticRegression Temp 0.062 ± 0.006 0.17 ± 0.039 0.058 ± 0.005 0.166 ± 0.042 0.833 ± 0.011 LogisticRegression Isotonic 0.002 ± 0.001 0.115 ± 0.001 0.002 ± 0.001 0.115 ± 0.001 0.754 ± 0.0 DecisionTree ERM 0.019 ± 0.001 0.039 ± 0.005 0.018 ± 0.002 0.038 ± 0.003 0.863 ± 0.001 DecisionTree HKRR 0.005 ± 0.001 0.033 ± 0.008 0.005 ± 0.001 0.032 ± 0.007 0.856 ± 0.001 DecisionTree HJZ 0.016 ± 0.003 0.042 ± 0.004 0.015 ± 0.002 0.036 ± 0.003 0.862 ± 0.0 DecisionTree Platt 0.02 ± 0.002 0.051 ± 0.007 0.018 ± 0.002 0.046 ± 0.007 0.861 ± 0.001 DecisionTree Temp 0.06 ± 0.002 0.08 ± 0.002 0.052 ± 0.001 0.069 ± 0.003 0.863 ± 0.002 DecisionTree Isotonic 0.006 ± 0.001 0.034 ± 0.004 0.009 ± 0.002 0.032 ± 0.003 0.863 ± 0.001 NaiveBayes ERM 0.134 ± 0.001 0.416 ± 0.006 0.126 ± 0.0 0.324 ± 0.003 0.808 ± 0.001 NaiveBayes HKRR 0.006 ± 0.001 0.032 ± 0.005 0.006 ± 0.001 0.03 ± 0.003 0.848 ± 0.001 NaiveBayes HJZ 0.008 ± 0.002 0.058 ± 0.009 0.011 ± 0.001 0.058 ± 0.009 0.754 ± 0.0 NaiveBayes Platt 0.008 ± 0.002 0.079 ± 0.002 0.012 ± 0.001 0.079 ± 0.001 0.754 ± 0.0 NaiveBayes Temp 0.185 ± 0.002 0.271 ± 0.004 0.184 ± 0.002 0.257 ± 0.003 0.809 ± 0.0 NaiveBayes Isotonic 0.002 ± 0.001 0.115 ± 0.001 0.002 ± 0.001 0.115 ± 0.001 0.754 ± 0.0
Figure 55: HMDA. Alternate groups.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ MLP ERM 0.017 ± 0.005 0.3 ± 0.057 0.021 ± 0.004 0.204 ± 0.03 0.866 ± 0.002 MLP HKRR 0.019 ± 0.002 0.311 ± 0.08 0.018 ± 0.002 0.217 ± 0.043 0.84 ± 0.002 MLP HJZ 0.025 ± 0.006 0.28 ± 0.049 0.024 ± 0.003 0.208 ± 0.031 0.864 ± 0.003 MLP Platt 0.02 ± 0.004 0.314 ± 0.048 0.023 ± 0.002 0.239 ± 0.039 0.863 ± 0.002 MLP Temp 0.063 ± 0.037 0.282 ± 0.061 0.058 ± 0.032 0.195 ± 0.038 0.864 ± 0.003 MLP Isotonic 0.025 ± 0.004 0.29 ± 0.047 0.025 ± 0.004 0.234 ± 0.027 0.864 ± 0.003 RandomForest ERM 0.019 ± 0.001 0.297 ± 0.038 0.021 ± 0.001 0.228 ± 0.018 0.863 ± 0.003 RandomForest HKRR 0.019 ± 0.002 0.311 ± 0.08 0.018 ± 0.002 0.217 ± 0.043 0.84 ± 0.002 RandomForest HJZ 0.016 ± 0.002 0.239 ± 0.036 0.019 ± 0.001 0.211 ± 0.015 0.861 ± 0.002 RandomForest Platt 0.019 ± 0.005 0.267 ± 0.03 0.021 ± 0.001 0.229 ± 0.02 0.86 ± 0.003 RandomForest Temp 0.041 ± 0.004 0.283 ± 0.057 0.039 ± 0.004 0.236 ± 0.014 0.861 ± 0.002 RandomForest Isotonic 0.015 ± 0.002 0.248 ± 0.019 0.017 ± 0.001 0.235 ± 0.012 0.862 ± 0.002 SVM ERM 0.143 ± 0.002 0.565 ± 0.0 0.072 ± 0.001 0.265 ± 0.0 0.857 ± 0.002 SVM HKRR 0.019 ± 0.002 0.311 ± 0.08 0.018 ± 0.002 0.217 ± 0.043 0.84 ± 0.002 SVM HJZ 0.027 ± 0.005 0.311 ± 0.037 0.026 ± 0.002 0.284 ± 0.027 0.826 ± 0.0 SVM Platt 0.14 ± 0.001 0.47 ± 0.051 0.07 ± 0.001 0.229 ± 0.02 0.86 ± 0.001 SVM Temp 0.117 ± 0.011 0.323 ± 0.022 0.116 ± 0.011 0.277 ± 0.018 0.847 ± 0.01 SVM Isotonic 0.048 ± 0.023 0.317 ± 0.076 0.048 ± 0.023 0.275 ± 0.052 0.847 ± 0.017 LogisticRegression ERM 0.022 ± 0.002 0.26 ± 0.02 0.022 ± 0.001 0.184 ± 0.009 0.866 ± 0.002 LogisticRegression HKRR 0.019 ± 0.002 0.311 ± 0.08 0.018 ± 0.002 0.217 ± 0.043 0.84 ± 0.002 LogisticRegression HJZ 0.017 ± 0.003 0.256 ± 0.064 0.02 ± 0.001 0.177 ± 0.028 0.863 ± 0.003 LogisticRegression Platt 0.02 ± 0.001 0.254 ± 0.032 0.022 ± 0.0 0.176 ± 0.031 0.862 ± 0.004 LogisticRegression Temp 0.102 ± 0.002 0.207 ± 0.045 0.094 ± 0.001 0.159 ± 0.025 0.862 ± 0.003 LogisticRegression Isotonic 0.016 ± 0.002 0.233 ± 0.03 0.019 ± 0.002 0.188 ± 0.037 0.861 ± 0.005 DecisionTree ERM 0.067 ± 0.004 0.328 ± 0.036 0.047 ± 0.004 0.189 ± 0.01 0.85 ± 0.006 DecisionTree HKRR 0.019 ± 0.002 0.311 ± 0.08 0.018 ± 0.002 0.217 ± 0.043 0.84 ± 0.002 DecisionTree HJZ 0.03 ± 0.007 0.286 ± 0.082 0.032 ± 0.007 0.225 ± 0.057 0.838 ± 0.006 DecisionTree Platt 0.106 ± 0.007 0.529 ± 0.064 0.059 ± 0.003 0.261 ± 0.03 0.836 ± 0.006 DecisionTree Temp 0.098 ± 0.003 0.3 ± 0.026 0.091 ± 0.002 0.232 ± 0.016 0.841 ± 0.005 DecisionTree Isotonic 0.055 ± 0.024 0.292 ± 0.039 0.042 ± 0.017 0.235 ± 0.046 0.836 ± 0.006 NaiveBayes ERM 0.277 ± 0.019 0.469 ± 0.031 0.164 ± 0.013 0.266 ± 0.025 0.714 ± 0.018 NaiveBayes HKRR 0.019 ± 0.002 0.311 ± 0.08 0.018 ± 0.002 0.217 ± 0.043 0.84 ± 0.002 NaiveBayes HJZ 0.027 ± 0.005 0.311 ± 0.037 0.026 ± 0.002 0.284 ± 0.027 0.826 ± 0.0 NaiveBayes Platt 0.268 ± 0.008 0.452 ± 0.012 0.164 ± 0.005 0.268 ± 0.01 0.719 ± 0.007 NaiveBayes Temp 0.298 ± 0.003 0.343 ± 0.006 0.276 ± 0.002 0.307 ± 0.004 0.719 ± 0.007 NaiveBayes Isotonic 0.017 ± 0.002 0.216 ± 0.06 0.019 ± 0.001 0.205 ± 0.055 0.829 ± 0.006
Figure 56: MEPS. Alternate groups.
Appendix H Results on Language and Image Datasets
H.1 Plots for All Multicalibration Algorithms
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 57: All multicalibration runs for image and language models. Note the small x-axis scale in some plots.
H.2 Result Tables for Image and Language Data
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ ViT ERM 0.021 ± 0.008 0.076 ± 0.011 0.022 ± 0.007 0.076 ± 0.011 0.965 ± 0.003 ViT HKRR 0.003 ± 0.0 0.018 ± 0.003 0.003 ± 0.0 0.018 ± 0.003 0.978 ± 0.0 ViT HJZ 0.006 ± 0.001 0.047 ± 0.001 0.007 ± 0.001 0.044 ± 0.003 0.973 ± 0.002 ViT Platt 0.014 ± 0.008 0.049 ± 0.014 0.016 ± 0.008 0.049 ± 0.014 0.973 ± 0.003 ViT Temp 0.025 ± 0.004 0.046 ± 0.01 0.019 ± 0.003 0.037 ± 0.008 0.973 ± 0.003 ViT Isotonic 0.001 ± 0.0 0.031 ± 0.007 0.002 ± 0.0 0.031 ± 0.007 0.977 ± 0.001 DenseNet-121 ERM 0.006 ± 0.003 0.047 ± 0.005 0.006 ± 0.002 0.047 ± 0.005 0.974 ± 0.001 DenseNet-121 HKRR 0.003 ± 0.002 0.018 ± 0.002 0.003 ± 0.002 0.018 ± 0.002 0.97 ± 0.003 DenseNet-121 HJZ 0.005 ± 0.001 0.056 ± 0.014 0.006 ± 0.001 0.055 ± 0.014 0.967 ± 0.003 DenseNet-121 Platt 0.006 ± 0.001 0.062 ± 0.015 0.007 ± 0.001 0.062 ± 0.015 0.971 ± 0.002 DenseNet-121 Temp 0.015 ± 0.002 0.052 ± 0.009 0.015 ± 0.002 0.05 ± 0.008 0.967 ± 0.003 DenseNet-121 Isotonic 0.002 ± 0.0 0.047 ± 0.006 0.003 ± 0.0 0.047 ± 0.006 0.972 ± 0.001
Figure 58: Camelyon17.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ ViT ERM 0.016 ± 0.006 0.069 ± 0.013 0.016 ± 0.006 0.068 ± 0.014 0.92 ± 0.005 ViT HKRR 0.008 ± 0.003 0.031 ± 0.003 0.008 ± 0.003 0.031 ± 0.003 0.926 ± 0.001 ViT HJZ 0.006 ± 0.001 0.038 ± 0.002 0.009 ± 0.0 0.037 ± 0.002 0.925 ± 0.001 ViT Platt 0.009 ± 0.002 0.047 ± 0.005 0.012 ± 0.002 0.047 ± 0.005 0.924 ± 0.001 ViT Temp 0.028 ± 0.008 0.072 ± 0.012 0.029 ± 0.009 0.07 ± 0.014 0.917 ± 0.006 ViT Isotonic 0.005 ± 0.001 0.057 ± 0.005 0.007 ± 0.001 0.057 ± 0.005 0.922 ± 0.001 ResNet-50 ERM 0.008 ± 0.001 0.028 ± 0.001 0.009 ± 0.001 0.028 ± 0.001 0.945 ± 0.0 ResNet-50 HKRR 0.006 ± 0.001 0.024 ± 0.004 0.006 ± 0.001 0.024 ± 0.004 0.934 ± 0.006 ResNet-50 HJZ 0.006 ± 0.001 0.032 ± 0.003 0.008 ± 0.0 0.033 ± 0.003 0.934 ± 0.007 ResNet-50 Platt 0.005 ± 0.001 0.037 ± 0.004 0.007 ± 0.0 0.037 ± 0.004 0.935 ± 0.006 ResNet-50 Temp 0.017 ± 0.006 0.046 ± 0.006 0.017 ± 0.006 0.045 ± 0.006 0.933 ± 0.007 ResNet-50 Isotonic 0.003 ± 0.001 0.051 ± 0.009 0.006 ± 0.001 0.051 ± 0.009 0.933 ± 0.007
Figure 59: CelebA.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ DistilBERT ERM 0.021 ± 0.001 0.065 ± 0.005 0.021 ± 0.001 0.06 ± 0.004 0.915 ± 0.001 DistilBERT HKRR 0.013 ± 0.0 0.047 ± 0.005 0.013 ± 0.0 0.043 ± 0.004 0.915 ± 0.001 DistilBERT HJZ 0.004 ± 0.001 0.043 ± 0.008 0.007 ± 0.001 0.043 ± 0.007 0.915 ± 0.001 DistilBERT Platt 0.004 ± 0.001 0.047 ± 0.008 0.007 ± 0.0 0.045 ± 0.007 0.915 ± 0.001 DistilBERT Temp 0.025 ± 0.005 0.044 ± 0.004 0.025 ± 0.005 0.044 ± 0.004 0.914 ± 0.001 DistilBERT Isotonic 0.002 ± 0.0 0.032 ± 0.006 0.005 ± 0.0 0.032 ± 0.006 0.916 ± 0.0
Figure 60: Civil Comments.
{adjustwidth}
-1in-1in Model ECE $\downarrow$ Max ECE $\downarrow$ smECE $\downarrow$ Max smECE $\downarrow$ Acc $\uparrow$ ResNet-56 ERM 0.039 ± 0.013 0.094 ± 0.009 0.039 ± 0.013 0.094 ± 0.009 0.867 ± 0.001 ResNet-56 HKRR 0.015 ± 0.001 0.059 ± 0.01 0.015 ± 0.001 0.047 ± 0.005 0.848 ± 0.004 ResNet-56 HJZ 0.013 ± 0.005 0.081 ± 0.012 0.014 ± 0.005 0.081 ± 0.012 0.863 ± 0.002 ResNet-56 Platt 0.009 ± 0.003 0.082 ± 0.01 0.01 ± 0.002 0.082 ± 0.01 0.863 ± 0.002 ResNet-56 Temp 0.024 ± 0.01 0.07 ± 0.003 0.024 ± 0.01 0.069 ± 0.003 0.863 ± 0.002 ResNet-56 Isotonic 0.005 ± 0.001 0.079 ± 0.009 0.007 ± 0.0 0.078 ± 0.008 0.863 ± 0.002
Figure 61: Amazon Polarity.